### v1

# Low-dimensional tori in Calogero–Moser–Sutherland systems

Abstract.
We give an explicit description of low-dimensional Liouville tori in Calogero–Moser–Sutherland systems.
1. Introduction
Recall that Calogero–Moser–Sutherland system [1, 6, 8] is a model of $n$ one-dimensional particles on a circle with trigonometric potential. The phase space of this model is $T^{*}H_{\rm reg}$ with coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$.
The symplectic structure is standard with Poisson brackets:
(1) $\{p_{i},p_{j}\}=0,\quad\{p_{i},q_{j}\}=\delta_{ij},\quad\{q_{i},q_{j}\}=0.$
Here we used coordinates $q_{j},\gamma_{j}=e^{iq_{j}}$.
The Hamiltonian of the system is
(2) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}(q_{i}-q_{j})}.$
This is an integrable system with higher integrals
(3) $H_{k}=\frac{1}{k}\mathrm{Tr}(x^{k}),$
where $x$ is given by formula (12). Function $H_{k}$ are invariant with respect to simultenious permutations of $p$ and $q$. Thus, effective phase space is an orbifold $T^{*}H_{reg}/S_{n}$, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
The main result of this paper is the following explicit description of all Liouville tori. Let $(x_{1},\dots,x_{n})$ be the eigenvalues of $x$ (given (12)). Then:
  • when $x_{i}-x_{i+1}>c$ for all $i$, the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-1)$-dimensional and with well-known angle coordinates (see section 2.3).
  • when $x_{k_{i}}-x_{k_{i}+1}=c$ for some $1\leq k_{1}<\dots<k_{\ell}\leq n$, then the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-\ell-1)$-dimensional (the generic case corresponds to $\ell=0$), with the angle variables described in section 3.
Generic $(n-1)$-dimensional tori are well known and completely described long time ago [7]. When $\ell=n-1$, the level surface is zero-dimensional, and is a fixed point of the multitime evolution (see also [LRS]).
$x_{2}-x_{3}$$x_{1}-x_{2}$cc2222-dim1111-dim1111-dim00-dim
Figure 1. Dimensions of Liouville tori over shifted principal Weyl chamber in the $x$-plane for $G=SU(3)$.
This paper was inspired and motivated by the work [5] on hybrid integrable systems and on the upcoming paper [2] on stratified integrable systems.
Similar stratification exists in many other models, for example, in the elliptic CM system and in integrable systems on moduli spaces of flat connections. We will address this in a separate publication. Superintegrable systems and, in particular, superintegrable systems of spin CMS type are studied in [2].
Acknowledgment
We are grateful to L. Feher for stimulating discussions and pointing out reference [3], to K. Jiang, Z. Chen, and H. Xiao for many discussions on stratified integrable systems.
The work of A. L. was supported by the Beijing Natural Science Foundation (IS24006). The work of N. R. was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the RF.
2. Action-angle variables in the Calogero–Moser–Sutherland system
2.1. The Calogero–Moser–Sutherland system for $SU(n)$.
Here we review well known facts about the Calogero–Moser (CM) system for $SU(n)$. This section is a well-known extension of [4] from $T^{*}\mathfrak{g}$ to $T^{*}G$. We included it for clarity of exposition. Elements of $\mathfrak{su}(n)$ and traceless $n\times n$ hermitian matrices and elements of $SU(n)$ are unimodular $n\times n$ unitary matrices.
The phase space for the Calogero–Moser system for $G=SU(n)$(and of its spin analogs) can be realized as the symplectic reduction of $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the Hamiltonian action of $G$ by conjugations. From now we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form. For $\mathfrak{su}(n)$ this means that both $\mathfrak{su}(n)$ and $\mathfrak{su}(n)^{*}$ are represented by traceless $n\times n$ matrices. If $\mathcal{O}$ is a coadjoint orbit in $\mathfrak{su}(n)$ of rank one, the corresponding Hamiltonian reduction is
(4) $\mathcal{S}(\mathcal{O})=\{(x,g)\in\mathfrak{su}(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G,$
In general, it is a stratified symplectic space. When $\mathcal{O}$ is a rank $1$ orbit, it is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n},$ where $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ and is a symplectic orbifold. Since the Hamiltonians are also defined on a smooth manifold $T^{*}H_{reg}$, it is convenient to work there. We will denote it $\widetilde{\mathcal{S}(\mathcal{O})}$.
A rank $1$ orbit in $\mathfrak{su}(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
$\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
modulo transformations $\varphi_{j}\mapsto e^{i\theta}\varphi_{j}$. It is clear that the second Casimir function 111 The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (5) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
$\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
The moment map brings the $SU(n)$-orbit through $(x,g)\in T^{*}G$ to
(6) $\mu=x-g^{-1}xg.$
From the construction of the phase space as a Hamiltonian reduction of $T^{*}SU(n)$ with respect to rank $1$ orbit the functions
(7) $H_{k}=\frac{1}{k}\mathrm{Tr}\big{(}x^{k}\big{)},\ \ k=2,\dots,n$
on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$ Poisson commute. This follows immediately from the symplectic structure on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$. As a consequence the Poisson commute on $\mathcal{S}(\mathcal{O})$. Functions $\{H_{k}\}_{k=2}^{n}$ are independent on $S(\mathcal{O})$ and because $\mathrm{dim}\,S(\mathcal{O})=2n-2$ they define a Liouville integrable system.
2.2. The diagonalization of $g$ and natural coordinates
Let us briefly recall the Hamiltonian reduction construction of Calogero-Moser system by Kazhdan, Kostant, and Sternberg [4], extended to Calogero-Moser-Sutherland. We can choose a representative of the $SU(n)$–orbit through $(x,g)$ in (4) such that $g$ is diagonal:
(8) $g=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n}).$
The equation (6) for diagonal $g$ becomes
(9) $x_{ij}(1-\gamma_{i}^{-1}\gamma_{j})=\overline{\varphi}_{i}\varphi_{j},\quad$ $i\neq j,$
(10) $0=|\varphi_{i}|^{2}-c,\quad\qquad\qquad$ $i=j.$
Thus $|\varphi_{i}|^{2}=c,\ c>0$.
Note that if $\gamma_{i}=\gamma_{j}$, the equation (9) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus $\gamma_{i}\neq\gamma_{j}$.
When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (8) is the Cartan subgroup of $SU(n)$ of unitary diagonal unimodular matrices. By the action of this subgroup on $x$ we can bring $\varphi_{j}$ to the form $\varphi_{j}=\sqrt{c}\,e^{i\theta}$. This gives the description of the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $(x_{ii},\gamma_{i})$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
For $j\neq k$ we have
(11) $x_{jk}=\frac{c}{1-\gamma_{j}^{-1}\gamma_{k}}.$
As it should be, we have $\bar{x}_{kj}=x_{jk}$.
Setting $p_{i}:=x_{ii}$ we have
(12) $x=\begin{pmatrix}p_{1}&\frac{c}{1-\gamma_{1}^{-1}\gamma_{2}}&\ldots&\frac{c}{1-\gamma_{1}^{-1}\gamma_{n}}\\ \frac{c}{1-\gamma_{2}^{-1}\gamma_{1}}&p_{2}&\ldots&\ldots\\ \ldots&\ldots&\ldots&\ldots\\ \frac{c}{1-\gamma_{n}^{-1}\gamma_{1}}&\ldots&\ldots&p_{n}\end{pmatrix}.$
The unimodularity implies $\sum_{i=1}^{n}p_{i}=0$. For the first two Hamiltonians we have:
(13) $H_{2}=\frac{1}{2}\mathrm{Tr}\big{(}x^{2}\big{)}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{\begin{subarray}{c}i<j\end{subarray}}^{n}\frac{c^{2}}{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})},$
(14) $H_{3}=\frac{1}{3}\mathrm{Tr}\big{(}x^{3}\big{)}=\frac{1}{3}\sum_{i=1}^{n}p_{i}^{3}+\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}\,p_{i}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{3}\sum_{i=1}^{n}p_{i}^{3}+\frac{1}{4}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}\,p_{i}}{{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})}}.$
Here $\gamma_{j}=\exp(iq_{j})$.
2.3. The diagonalization of $x$ and the action-angle variables
Geometrically, any integrable system is a projection from a $2n$-dimensional symplectic manifold to an $n$-dimensional space, such that the generic fiber is Lagrangian.
For Calogero–Moser systems we have the projection
$\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(x,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
Here $x$ is given by (12) as the function of $p$ and $\gamma$, and for $\mathfrak{su}(n)$. $W=S_{n}$, and $[x]$ is a collection of eigenvalues of $x$. The quotient space $\mathfrak{h}^{*}/W$ can be modeled as:
$\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}.$
The main observation of this paper is that the image of this mapping is a shifted cone in $\mathbb{R}^{n-1}$. Preimages of strata of this cone stratify $T^{*}H_{\mathrm{reg}}$. For CMS model we describe the strata later in this section.
2.3.1. The generic stratum
Any $x\in\mathfrak{su}(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$. let us choose a representative $(x,g)$ in our phase space where $x$ is diagonal:
(15) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset\mathfrak{su}(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
The Hamiltonians, evaluated in terms of eigenvalues of $x$ are simply symmetric power functions:
$H_{k}=\sum_{j=1}^{n}x_{j}^{k}.$
Theorem 1.
Solutions to the momentum map equation
(16) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\cdot 1.$
can be described in the following way. Let’s $\lambda=g\bar{\psi}$, then the following holds
1) The following conditions are equivalen
  • $x_{i}-x_{i+1}\geq c$ for all $i=1,\ldots,n-1$,
  • $|\psi_{i}|^{2}\geq 0$ for all $i=2,\ldots,n$,
  • $|\lambda_{i}|^{2}\geq 0$ for all $i=1,\ldots,n-1$.
Moreover, $|\psi_{1}|^{2}>0$ and $|\lambda_{n}|^{2}>0$.
2) For any $n>k\geq 1$, the following occurs only at the same time
  • $x_{k}-x_{k+1}=c$,
  • $|\psi_{k+1}|^{2}=0$,
  • $|\lambda_{k}|^{2}=0$.
3) If $x_{i}-x_{i+1}>c$ for $i=1,\ldots,n-1$, then
(17) $|\psi_{i}|^{2}=c\prod_{j=1,j\neq i}^{n}\frac{x_{i}-x_{j}+c}{x_{i}-x_{j}},\qquad\text{and }\qquad|\lambda_{i}|^{2}=c\prod_{j=1,j\neq i}^{n}\frac{x_{j}-x_{i}+c}{x_{j}-x_{i}}.$
4) For $x_{i}-x_{i+1}>c$ with $i=1,\ldots,n-1$ matrix $g$ has form
(18) $g_{ij}=\frac{\lambda_{i}\,\psi_{j}}{x_{j}-x_{i}+c},$
Proof.
1) The equation for the moment map can be written as
(19) $g^{-1}xg=x+c\cdot 1-\bar{\psi}\psi^{t},$
Following lemma 3.4 from [3], let’s consider the equality of characteristic polynomials of both parts of the equation (31)
(20) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c\cdot 1-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
Here we take into account that for diagonal matrix $A$
(21) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{adj}u,$
where $A^{adj}$ is the adjugate matrix to $A$, which defined by $A^{adj}A=\mathrm{det}A\cdot 1$. This is a well-known fact which follows from Lemma 1 in the appendix A.
Equation (20) at the point $z=x_{i}+c$ is
(22) $\prod_{j=1}^{n}(x_{i}-x_{j}+c)=|\psi_{i}|^{2}\prod_{j\neq i}(x_{i}-x_{j}).$
It follows from the equation (22) for $i=1$ that $|\psi_{1}|$ is strictly positive
(23) $|\psi_{1}|^{2}=c\prod_{j>1}\frac{x_{1}-x_{j}+c}{x_{1}-x_{j}}>0.$
In equation (22) for $i=2$
(24) $c\ \frac{(x_{2}-x_{1}+c)}{x_{2}-x_{1}}\prod_{j>2}\frac{(x_{2}-x_{j}+c)}{(x_{2}-x_{j})}=|\psi_{2}|^{2}$
the product over $j>2$ is strictly positive. So we have
(25) $|\psi_{2}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{1}\geq x_{2}+c,$
where equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
Continuing this reasoning, we get the statement of the proposition. So, we have
(26) $|\psi_{k}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
where equality $x_{k-1}=x_{k}+c$ is achieved if $|\psi_{k}|^{2}=0$.
The equation for the moment map can be rewritten for $\lambda$ as
(27) $gxg^{-1}=x-c\cdot 1+\lambda\bar{\lambda}^{t},$
Equality of characteristic polynomials gives
(28) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c\cdot 1+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
Equation (28) at the point $z=x_{i}-c$ is
(29) $\prod_{j=1}^{n}(x_{j}-x_{i}+c)=|\lambda_{i}|^{2}\prod_{j\neq i}(x_{j}-x_{i}).$
Using a similar argument as for equation (22), we get $|\lambda_{n}|^{2}>0$ and
(30) $|\lambda_{k-1}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
where equality $x_{k-1}=x_{k}+c$ is achieved if $|\lambda_{k-1}|^{2}=0$.
2) This part of the statement directly follows from (22) and (29).
3) After the multiplication of the momentum map (31) by $g$ the equation for the moment map becomes
(31) $gx+cg-xg=\lambda\psi^{t},$
$(i,j)$ whose element has the form
(32) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
which fix the form of the thirt part of the statement
which implies (18).
Remark 1.
If $x_{i}>x_{i+1}+c$, the stabilizer of $x$ is the Cartan subgroup in $SU(n)$, i.e., unitary diagonal matrices. Acting by this stabilizer on $\psi$, we can bring it to the form $\psi_{i}=|\psi_{i}|$ for all $i$. After this, we have:
(33) $g=\alpha X,\qquad X_{ij}=\frac{|\lambda_{i}||\psi_{j}|}{x_{j}-x_{i}+c},\qquad\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,$
where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. with $\mathrm{det}\,X=1$ according to
The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
Remark 2.
We can multiply $g$ by a central element $\omega\in Z(SU(n))$:
(34) $\omega=\exp\left(\frac{2\pi i}{n}k\right)\cdot 1_{n},\qquad\text{ where }k=0,\ldots,n-1.$
If $\gamma_{j}$ are eigenvalues of the matrix $g$, such multiplication changes $\gamma_{j}\mapsto\omega\gamma_{j}$. Eigenvalues of $g$ are exponents of coordinates of Calogero–Moser–Sutherland model $\gamma_{j}=\exp(iq_{j})$. Thus the multiplication of $g$ by $\omega$ shifts coordinates as
[…]
Thus, we described coordinates on the generic (maximal) stratum: real coordinates $\{x_{i}\}$ and $\{\mathrm{arg}\,\alpha_{i}\}$ with satisfying constrains $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\prod_{i=1}^{n}\alpha_{i}=1$, $|\alpha_{i}|=1$.
Now let us describe all strata and let us show that $(x,\alpha)$ are the action-angle variables.
Define $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
(36) $x_{i}=\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}.$
Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\log\alpha_{k}$ for $j=2,\ldots,n$. Then,
(37) $\alpha_{k}=e^{i(\theta_{k+1}-\theta_{k})},\qquad k=1,\ldots n.$
Here we assume that $\theta_{1}=\theta_{n+1}=0$.
Thus, we get coordinates on the generic (maximal) stratum: real coordinates $(y_{2}\ldots,y_{n})$ satisfying constrains $y_{i}>c$ and $(\theta_{2}\ldots,\theta_{n})$. In section 4 we show that $(y,\theta)$ are the action-angle variables.
3. Low–dimensional tori
The moment map equation (6) when $x$ is diagonal becomes
(38) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
where $\lambda=g\bar{\psi}$.
Proposition 1.
If for some $k$ we have $\psi_{k}=0$, then in the $k$-th column and $(k-1)$-th row of the martrix $g$ only nonzero element is $g_{k-1,k}\neq 0$ and $|g_{k-1,k}|=1$.
Proof.
Consider equation (38) for $j=k$
(39) $(x_{k}-x_{i}+c)g_{ik}=0.$
If for all $i$ elements $g_{ik}=0$ then $g$ not invertible. Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$. For $j\neq k-1$ $x_{j}-x_{k}+c\neq 0$ and therefore $g_{jk}=0$. This means that $k$-th column has only one non-zero element $g_{k-1,k}$.
Unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the $(k-1)$-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$. ∎
Assume that $\psi_{j}\neq 0$ only for $j=1,j_{1},\dots,j_{s}$ where $j_{1}<j_{2}<\dots<j_{s}$. For convenience, define $j_{0}=1$.
Define the matrix $\widehat{g}$ as $P\widehat{g}=g$, where $P$ is the matrix of cyclic permutation with sign
(40) $P=\begin{pmatrix}0&-1&0&\ldots&0\\ 0&0&-1&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&-1\\ 1&0&0&\ldots&0\\ \end{pmatrix}.$
Proposition 1 implies
$\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
if neither $i$ no $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
The moment map equations give the following identity
$x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c$
where $x_{j}^{P}=x_{j-1}$ with $x_{0}=x_{n}$, and equations
(41) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}$
Define
$\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
Theorem 2.
The following holds:
  • $\widehat{\lambda}_{i}\neq 0,\lambda_{i}$, for $i\neq j_{a}$, $a=0,\dots,s$.
  • (42) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})},$
  • (43) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}$
  • By the action of the Cartan part of the stabilizer of $x$, we can bring $\widehat{g}$ to the form
    (44) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}$
Proof.
Write the moment map equation as
$\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
For the characteristic polynomial of the right side, using lemma 1 from the appendix A we obtain
$\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
Taking $z=x_{j_{a}}+c$ we obtain
$\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
which proves the formula (42) for $|\psi_{j_{a}}|^{2}$.
Similarly, writing the moment map equation as
$\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
and using lemma 1 gives the proof of the formula (43) for $|\widehat{\lambda}_{j_{a}}|^{2}$.
Finally, the moment map equations (41) give
$\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives (44). ∎
Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer.
Remark 3.
We can write $g$ as the product
(45) $g=\alpha\,P\,Y,$
where $P$ is the cyclic permutation matrix (40).
Let $\beta$ be an element of the stabilizer $x$, i.e. a diagonal unitary matrix.
We use transformations
(46) $g\to\beta g\beta^{-1}$
with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real.
Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
$\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$.
Thus, we can fix the section of the action of the stabilizer group $H\subset SU(n)$:
(47) $g=\alpha\,P\,Y,$
where
$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\}$
(48) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad i\notin\{j_{0},\ldots,j_{s}\}$
(49) $Y_{j_{a}j_{b}}$ $=(-1)^{1-\delta_{a,0}}\frac{|\widehat{\lambda}_{j_{a}}||\psi_{j_{b}}|}{x_{j_{b}}-x_{j_{a}}^{P}+c},$
(50) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a},$
and $P$ is defined (40).
Note that with this choice, we have
(51) $\mathrm{det}\,\alpha=\mathrm{det}\,P=\mathrm{det}\,Y=1$
and real matrix $Y$ is othogonal
$YY^{t}=I.$
Define $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1},$ then due to (36)
(52) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c).$
Thus, we described coordinates $(y_{j_{a}},\alpha_{j_{a}})$ on $s$-dimention strata: $y_{j_{a}}>c$ and $\{\alpha_{j_{a}}\}_{a=0}^{s}$ with constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$.
Example 1.
(The fixed point). In the case when all $\psi_{i}=0$ for $i\geq 2$, we have
  • $|\psi_{1}|^{2}=|\hat{\lambda}_{1}|^{2}=nc$.
  • $x_{i}=\frac{n-1}{2}c-(i-1)c$ for $1=2,\ldots,n$.
(53) $g=P.$
In terms of original CMS coordinates we have $q_{j}=\frac{2\pi j}{n}$. The action of the center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}\mod(2\pi)$.
Example 2.
(One-domensional tori). Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ when $i\neq 1,j$.
  • $|\psi_{1}|^{2}+|\psi_{j}|^{2}=|\hat{\lambda}_{1}|^{2}+|\hat{\lambda}_{j}|^{2}=nc$.
  • For $y=x_{j-1}-x_{j}$, $y>c$; $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$, and with $\sum_{i=1}^{n}x_{i}=0$:
    (54) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c),$
    where $\delta_{i\geq j}$ is a step function.
  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j},1,\ldots,1\right)$.
(55) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c}$
and
(56) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix},$
where
(57) $r^{2}=\frac{(j-1)c\,(n-j+1)c}{(y+(j-2)c)(y+(n-j)c)},\qquad q^{2}=\frac{(y-c)(y+(n-1)c)}{(y+(j-2)c)(y+(n-j)c)}.$
Note that $r^{2}+q^{2}=1$. It is clear that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.
Remark 4.
Theorem 2 can be obtained from theorem 1 directly using the appropriate limit. In particular, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$ ($j_{0}=1$)
(58) $|\psi_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a},$
(59) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{for }i=j_{a},$
where in the LHS $|\psi_{i}|^{2}$ from (17) and in the RHS $|\psi_{j_{a}}|^{2}$ from (42), and
(60) $|\lambda_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a}-1,$
(61) $|\lambda_{i}|^{2}$ $\to|\widehat{\lambda}_{j_{a}}|^{2},\qquad\text{for }i=j_{a}-1\ \mathrm{mod}\ n.$
4. Action-angle variables
The cotangent bundle $T^{*}G$ carries the natural symplectic form
(62) $\Omega=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right).$
Here we used the trivialization $T^{*}G\simeq\mathfrak{g}^{*}\times G$ by right translation. The identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$ is given by the Killing form. We also assume that $G$ is a compact matrix group.
This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. After the trivialization and identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$, it acts as
$h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
The symplectic form (62) restricted to $\mu^{-1}(\mathcal{O})$ gives a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$. We can model $S(\mathcal{O})$ as the space of pairs $(x,g)$, up to the action $H\subset SU(n)$, where $x=(x_{1}\geq x_{2}\geq\ldots\geq x_{n})$, $\sum_{i=1}^{n}x_{i}=0$, is diagonalized. On the big stratum $g=\alpha X$ (33), on others strata $g=\alpha PY$, see remark 3.
Now, let us describe the restriction $\Omega_{\mathcal{O}}$ of the symplectic form (62) in terms of $(y,\theta)$ coordinates.
Proposition 2.
On the generic stratum
(63) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j}.$
where $x_{i}$ are eigenvalues of $x$ and $\alpha_{j}=\mathrm{exp}(i\eta_{j})$.
Proof.
Using the representation (33) for $\Omega$ form we get
(64) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha X)\,X^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right).$
The first term vanishes:
(65) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0,$
where we take into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$.
Thus
(66) $\Omega_{\mathcal{O}}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right)=i\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\alpha^{-1}_{i}\mathrm{d}\alpha_{i}=-\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\eta_{i}.$
Taking into account constraints $\sum_{j=1}^{n}\mathrm{d}x_{j}=0$ and $\sum_{j=1}^{n}\mathrm{d}\eta_{j}=0$ we can write $\Omega_{\mathcal{O}}$ as
(67) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j}.$
where $y_{j}=x_{j-1}-x_{j}$ and $\theta_{j}=\sum_{i=1}^{j-1}\eta_{i}$. Thus, on the generic stratum $(y_{j},\theta_{j})$ are Darboux action-angle coordinates.
Proposition 3.
On a $2s$-dimensional stratum of $S(\mathcal{O})$ the symplectic form $\Omega_{\mathcal{O}}$ becomes
(68) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}}.$
where $y_{j_{a}}=x_{j_{a}-1}-x_{j_{a}}$ and $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$, with $\alpha_{j_{b}}=e^{i\eta_{j_{b}}}$.
Proof.
Using representation (47) for $\Omega$ form we get
(69) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha PY)\,Y^{-1}P^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}Y\,Y^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x^{P}\,\mathrm{d}\alpha\,\alpha^{-1}\right),$
where $x^{P}=P^{-1}xP$.
Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrains $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies.
(70) $\eta_{j_{0}}\equiv\eta_{1}=-\sum_{a=1}^{s}\eta_{j_{a}}.$
Expression $\mathrm{Tr}\left(x^{P}\,\mathrm{d}Y\,Y^{-1}\right)=0$ for the same reasons as in equation (65) and $\Omega_{\mathcal{O}}$ becomes
(71) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=\\ -\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=\\ -\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
Here we used (70). In term of $\theta_{j_{a}}=\sum_{a=b}^{s}\eta_{j_{b}}$ this gives (68). ∎
Remark 5.
Note that $\theta$’s defined in proposition 3 differ slightly from $\theta$’s defined in proposition 2. This difference arises due to the introduction of a shift matrix $P$ (40).
The coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux action-angle coordinates on this $s$-dimensional stratum.
5. The multitime dynamics of Calogero–Moser–Sutherland system
5.1. The multi-time dynamics on $T^{*}G$.
Let $v_{k}$ be the Hamiltonian vector field on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ generated by $H_{k}=\frac{1}{k}\operatorname{tr}(x^{k})$, i.e. $i_{v_{k}}\Omega=-dH_{k}$, where $\Omega$ is the symplectic form (62) on $T^{*}G$. We have
(72) $v_{k}=-i\sum_{i,j=1}^{n}\left(x^{k-1}g\right)_{ij}\frac{\partial}{\partial g_{ij}},$
Because Hamiltonians $H_{k}$ Poisson commute on $T^{*}G$, vector fields $v_{k}$ also commute. Their multi-time flows are solutions to the system of differential equations:
(73) $\frac{\mathrm{d}x({\bf t})}{\mathrm{d}t_{k}}$ $=0,$
(74) $\frac{\mathrm{d}g({\bf t})}{\mathrm{d}t_{k}}$ $=ix({\bf t})^{k-1}g({\bf t}).$
The initial value problem $x(0)=x,\,g(0)=g$ for these equations can be solved explicitly:
(75) $x({\bf t})=x,\qquad g({\bf t})=e^{i\sum_{k=1}^{n}t_{k}\,x^{k-1}}g.$
This dynamics preserves Hamiltonians $H_{k}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
5.2. The multi-time dynamics on the generic stratum.
In coordinates $(y,\theta)$ the symplectic form on $S(\mathcal{O})$ obtained by the Hamiltonian reduction from (62) is
(76) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j},$
which means Poisson brackets between coordinate function $\{y_{i}\}$ and $\{\theta_{i}\}$ are
$\{y_{i},y_{j}\}=0,\qquad\{\theta_{i},\theta_{j}\}=0,\qquad\{\theta_{i},y_{j}\}=\delta_{ij}.$
Any $G$-invariant function $H$ on $T^{*}G$ defines a $G$-invariant function on $\mu^{-1}(\mathcal{O})\subset T^{G}$ and a function $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ which we denote by same letter $H$. Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
(77) $v_{k}=\sum_{i,j}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j}}\frac{\partial}{\partial\theta_{j}},\qquad\text{ for }k=2,\ldots,n.$
Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
(78) $v_{k}=\sum_{j=2}^{n}u^{k}_{j}\frac{\partial}{\partial\theta_{j}},$
where
(79) $u^{k}_{j}=\sum_{i=1}^{n}\left(\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}\right)^{k-1}\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right).$
Flow lines of these vector fields are linear in coordinates $\theta_{l}$. The multitime flow line with $y_{j}(0)=y_{j}$, $\theta_{j}(0)=\theta_{j}$ is
(80) $y_{j}({\bf t})=y_{j},\qquad\theta_{j}({\bf t})=\theta_{j}-\sum_{k=1}^{n-1}u^{k}_{j}\,t_{k},\qquad\text{ for }k=2,\ldots,n.$
From this we have:
(81) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})Y,$
where $\alpha_{k}(\mathbf{t})=\exp\left(i(\theta_{k+1}(\mathbf{t})-\theta_{k}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{n+1}(\mathbf{t})=0$ and $Y$ is defined in (33).
The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
5.3. The multi-time dynamics on low-dimensional strata.
When $\psi_{i}=0$ except for $i=j_{0},j_{1},\ldots,j_{s}$ (here $j_{0}=1$), the matrix $g$ can be written as
(82) $g=\alpha PY.$
In coordinates $(y,\theta)$ the symplectic form on the correspond stratum of $S(\mathcal{O})$, see remark 3, is
(83) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
which implies
$\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
(84) $v_{k}=\sum_{i=1}^{n}\sum_{a=1}^{s}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
Here we can choose $s$ independent vector fields. In particular, we can choose the first $s$ vectors.
Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
(85) $v_{k}=\sum_{a=2}^{s}u^{k}_{a}\frac{\partial}{\partial\theta_{j_{a}}},$
where
(86) $u^{k}_{a}=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c)\right)^{k-1}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right).$
The solution to the initial value problem $y_{j_{a}}(0)=y_{j_{a}},\,\theta_{j_{a}}(0)=\theta_{j_{a}}$ for these equations is
(87) $y_{j_{a}}({\bf t})=y_{j_{a}},\qquad\theta_{j_{a}}({\bf t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{k}_{a}\,t_{k},\qquad\text{ for }a=1,\ldots,s.$
So,
(88) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})P\,Y,$
where $\alpha_{j_{a}}(\mathbf{t})=\exp\left(i(\theta_{j_{a}}(\mathbf{t})-\theta_{j_{a+1}}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{j_{s+1}}(\mathbf{t})=0$, $Y$ is defined in (48) and $\alpha$ is defined in (50).
The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
Example 3.
Zero-dimensional stratum.
On this stratum, particles are in equilibrium. Their position are terminated by eigenvalues of (53)
(89) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{n}=0.$
Example 4.
One-dimensional strata.
There are $(n-1)$ such strata, labeled by $j=2,\ldots,n$, with coordinates $(y_{j},\theta_{j})$ defined in example 2.
Positions of particles are eigenvalues of $g(t)$ and thus are roots of the characteristic equation
(90) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{j-1}q\beta^{-1}\lambda^{n-j+1}+(-1)^{n-j+1}q\beta\lambda^{j-1}+(-1)^{n}=0,$
where $\beta=\beta(t)=\alpha_{1}^{-1}=\alpha_{j}=\exp\left(i\theta_{j}(t)\right)=\exp(i\,(\theta_{j}-u\,t))$ and $10<q<$ 1 defined in example 2, and
(91) $u=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y_{j}-c)\right)\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)=\\ (n-j+1)(j-1)\left(\frac{c}{2}+\frac{y-c}{n}\right).$
Appendix A Lemma about determinant of sum of matrix plus rank $1$ matrix
Lemma 1.
Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a $1$-dimensional projector. Then
(92) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{\beta}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z)$
Proof.
We have
(93) $\det(D-z-P)$ $=\exp\left(\mathrm{Tr}\ln(D-P-z)\right)$
(94) $=\exp\left(\mathrm{Tr}\ln(D-z)+\operatorname{Tr}\ln\left(1-(D-z)^{-1}P\right)\right)$
(95) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)$
Now
(96) $\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}$ $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{n}}-z)^{-1}P_{a_{1}a_{2}}\cdots P_{a_{k}a_{1}}$
(97) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}$
From here:
$\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}$
Together with (93) this implies (92). ∎

### v2

@@
-Low-dimensional tori in Calogero–Moser–Sutherland systems
+Low–dimensional tori in Calogero–Moser–Sutherland systems
 Abstract.
-We give an explicit description of low-dimensional Liouville tori in Calogero–Moser–Sutherland systems.
+The main result of this paper is an explicit description of the stratification of the phase space of Calogero-Moser-Sutherland integrable systems corresponding to Lie groups $SU(n)$.
+Each stratum is isomorphic to $\mathbb{T}^{s}\times\mathbb{R}^{s}$ where $s=0,1,\dots,n-1$.
+Tori in such strata are Liouville tori.
+The zero-dimensional stratum is the fixed point of the multi-time dynamics of the model.
 1. Introduction
-Recall that Calogero–Moser–Sutherland system [1, 6, 8] is a model of $n$ one-dimensional particles on a circle with trigonometric potential.
-The phase space of this model is $T^{*}H_{\rm reg}$ with coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$.
-The symplectic structure is standard with Poisson brackets:
+Recall that Calogero–Moser–Sutherland (CMS) system [1, 6, 8] is a model of $n$ one–dimensional particles on a circle.
+The phase space of this model is $T^{*}H_{\rm reg}$ where $H_{\rm reg}$ is the regular part of the Cartan subgroup of $SU(n)$, i.e.
+diagonal unitary matrices with distinct eigenvalues.
+The space $T^{*}H_{\rm reg}$ has natural coordinates $(p_{i},\gamma_{i})$, $p_{i}\in\mathbb{R}$, $\lvert\gamma_{i}\rvert=1$, $\sum_{i=1}^{n}p_{i}=0$, $\prod_{i=1}^{n}\gamma_{i}=1$, $\gamma_{i}\neq\gamma_{j}$.
+The standard symplectic structure on $T^{*}H_{\rm reg}$ gives the following Poisson brackets:
 (2) $\{p_{i},p_{j}\}=0,\quad\{p_{i},q_{j}\}=\delta_{ij},\quad\{q_{i},q_{j}\}=0.$
@@
 The Hamiltonian of the system is
-(2) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}(q_{i}-q_{j})}.$
+(3) $H_{2}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{i\neq j}^{n}\frac{c^{2}}{\sin^{2}\left(\tfrac{q_{i}-q_{j}}{2}\right)}.$
 This is an integrable system with higher integrals
 (4) $H_{k}=\frac{1}{k}\mathrm{Tr}(x^{k}),$
-where $x$ is given by formula (12).
-Function $H_{k}$ are invariant with respect to simultenious permutations of $p$ and $q$.
-Thus, effective phase space is an orbifold $T^{*}H_{reg}/S_{n}$, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
-The main result of this paper is the following explicit description of all Liouville tori. Let $(x_{1},\dots,x_{n})$ be the eigenvalues of $x$ (given (12)). Then:
-  • when $x_{i}-x_{i+1}>c$ for all $i$, the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-1)$-dimensional and with well-known angle coordinates (see section 2.3).
-  • when $x_{k_{i}}-x_{k_{i}+1}=c$ for some $1\leq k_{1}<\dots<k_{\ell}\leq n$, then the level surface of $\left\{H_{k}\right\}_{k=2}^{n}$ is $(n-\ell-1)$-dimensional (the generic case corresponds to $\ell=0$), with the angle variables described in section 3.
-Generic $(n-1)$-dimensional tori are well known and completely described long time ago [7].
-When $\ell=n-1$, the level surface is zero-dimensional, and is a fixed point of the multitime evolution (see also [LRS]).
-$x_{2}-x_{3}$$x_{1}-x_{2}$cc2222-dim1111-dim1111-dim00-dim
-Figure 1. Dimensions of Liouville tori over shifted principal Weyl chamber in the $x$-plane for $G=SU(3)$.
-This paper was inspired and motivated by the work [5] on hybrid integrable systems and on the upcoming paper [2] on stratified integrable systems.
-Similar stratification exists in many other models, for example, in the elliptic CM system and in integrable systems on moduli spaces of flat connections.
+where the matrix $x$ as a function of $p$ and $q$ is given by formula (18).
+The matrix $x$ is known as the Lax matrix for the CMS system.
+Functions $\{H_{k}\}_{k=2}^{n}$ are independent [^1] They are independent in the usual sense: $dH_{2}\wedge\dots\wedge dH_{n}$ is not identically zero.
+and Poisson commute.
+This follows immediately from the Hamiltonian reduction construction [4].
+Functions $H_{k}$ are invariant with respect to simultaneous permutations of $p$ and $q$.
+Thus, the effective phase space is a manifold $S(\mathcal{O})=T^{*}H_{reg}/S_{n}$ [^2] The notation $S(\mathcal{O})$ for the phase space is natural from the point of view of the Hamiltonian reduction, see section 2.1, where $S_{n}$ acts on $\mathfrak{h}^{*}\times H_{reg}$ diagonally.
+Geometrically, the CMS system is a Lagrangian fibration $\pi:S(\mathcal{O})\to\mathfrak{h}/S_{n}$.
+Here $\mathfrak{h}\simeq\mathbb{R}^{n-1}$ is the Cartan subalgebra of $su(n)$ and $\pi([(x,g)])=[x]$, where $[(x,g)]$ and $[x]$ are conjugacy classes of $(x,g)$ and of $x$ respectively, see section 2.1 for details.
+One of the main results of this paper is an explicit description of $B(\mathcal{O})=Im(\pi)$ as a stratified space which is the shifted principal Weyl chamber for $su(n)$:
+(5) $B(\mathcal{O})=\sqcup_{\{j\}}B(\mathcal{O})_{\{j\}}.$
+Here $\{j\}=\{j_{0},j_{1},\dots,j_{s}\}$, where $j_{0}=1<j_{1}<\dots<j_{s}$, and $0\leq s\leq n-1$.
+The stratum $B(\mathcal{O})_{\{j\}}$ is isomorphic to $\mathbb{R}^{s}$.
+On this stratum $x_{i}-x_{i+1}=c$ if $i$ is not in the set $\{j\}$ and $x_{i}-x_{i+1}>c$ if $i\in\{j\}$.
+The maximum stratum $B(\mathcal{O})_{\{1,2,\dots,n\}}$ has dimension n1n-1 and is the interior of the shifted principal Weyl chamber, where $x_{i}-x_{i+1}>c$ for all $i=1,\dots,n-1$.
+Other strata are its walls.
+As a consequence the phase space $S(\mathcal{O})$ gets a stratification by the preimages of $B(\mathcal{O})_{\{j\}}$.
+(6) $S(\mathcal{O})=\sqcup_{\{j\}}S(\mathcal{O})_{\{j\}},$
+where $S(\mathcal{O})_{\{j\}}=\pi^{-1}(B(\mathcal{O})_{\{j\}})$. In section 3.3 we prove that
+(7) $S(\mathcal{O})_{\{j\}}\simeq\mathbb{T}^{s}\times\mathbb{R}^{s}.$
+The maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ is subject to the Arnold–Liouville Theorem and its structure is well known, see for example [7].
+The action-angle variables on this stratum are described in section 2.3.
+Action-angle variables on low–dimensional strata are described in section 3.3.
+The minimal stratum $S(\mathcal{O})_{\{1\}}$ is the fixed point of the multi-time Hamiltonian flow generated by Poisson commuting Hamiltonians $H_{k}$ (see [5]).
+$x_{2}-x_{3}$$x_{1}-x_{2}$cc22-dim11-dim11-dim0-dim
+Figure 1. The shifted principal Weyl chamber for $G=SU(3)$ and dimensions of Liouville tori over the corresponding strata.
+This paper is closely related to the work [5] on hybrid integrable systems and to the upcoming paper [2] on stratified super-integrable systems.
+Similar stratification exists in many other models, for example, in the elliptic CMS system and integrable systems on moduli spaces of flat connections.
 We will address this in a separate publication.
-Superintegrable systems and, in particular, superintegrable systems of spin CMS type are studied in [2].
+We also do not discuss here the algebro-geometric setting.
 Acknowledgment
@@
 L.
-was supported by the Beijing Natural Science Foundation (IS24006).
+was supported by the Beijing Natural Science Foundation (IS24006) and Beijing Talent Program.
 The work of N.
 R.
-was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the RF.
-2. Action-angle variables in the Calogero–Moser–Sutherland system
-2.1. The Calogero–Moser–Sutherland system for $SU(n)$.
-Here we review well known facts about the Calogero–Moser (CM) system for $SU(n)$.
-This section is a well-known extension of [4] from $T^{*}\mathfrak{g}$ to $T^{*}G$.
-We included it for clarity of exposition.
-Elements of $\mathfrak{su}(n)$ and traceless $n\times n$ hermitian matrices and elements of $SU(n)$ are unimodular $n\times n$ unitary matrices.
-The phase space for the Calogero–Moser system for $G=SU(n)$(and of its spin analogs) can be realized as the symplectic reduction of $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the Hamiltonian action of $G$ by conjugations.
-From now we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form.
-For $\mathfrak{su}(n)$ this means that both $\mathfrak{su}(n)$ and $\mathfrak{su}(n)^{*}$ are represented by traceless $n\times n$ matrices.
-If $\mathcal{O}$ is a coadjoint orbit in $\mathfrak{su}(n)$ of rank one, the corresponding Hamiltonian reduction is
-(4) $\mathcal{S}(\mathcal{O})=\{(x,g)\in\mathfrak{su}(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G,$
-In general, it is a stratified symplectic space.
-When $\mathcal{O}$ is a rank $1$ orbit, it is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n},$ where $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ and is a symplectic orbifold.
-Since the Hamiltonians are also defined on a smooth manifold $T^{*}H_{reg}$, it is convenient to work there.
-We will denote it $\widetilde{\mathcal{S}(\mathcal{O})}$.
-A rank $1$ orbit in $\mathfrak{su}(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
-$\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
+was supported by the Collaboration Grant ”Categorical Symmetries” from the Simons Foundation, by the Changjiang fund, and by the project 075-15-2024-631 funded by the Ministry of Science and Higher Education of the Russian Federation.
+2. The Calogero–Moser–Sutherland system
+2.1. The Hamiltonian reduction
+Here we review some facts about the CMS system for $SU(n)$.
+From now on, $G$ is always $SU(n)$.
+This section is included for completeness of the exposition.
+It is a well-known $T^{*}G$ version of the Kazhdan-Kostant-Sternberg construction [4] of the Calogero-Moser system by the Hamiltonian reduction of $T^{*}\mathfrak{g}$.
+Here and below $\mathfrak{g}=su(n)$. We treat elements of $su(n)$ as traceless $n\times n$ Hermitian matrices, and elements of $SU(n)$ as $n\times n$ unimodular unitary matrices.
+The phase space for the CMS system for $G=SU(n)$ can be realized as the Hamiltonian reduction of $T^{*}G\simeq\mathfrak{g^{*}}\times G$ with respect to the Hamiltonian action of $G$ by conjugations.
+From now on, we identify $\mathfrak{g}\simeq\mathfrak{g}^{*}$ via the Killing form.
+For $su(n)$ this means that both $su(n)$ and $su(n)^{*}$ are represented by traceless Hermitian $n\times n$ matrices.
+The Hamiltonian reduction for the adjoint action of $SU(n)$ on $T^{*}SU(n)$ via the coajoint orbit $\mathcal{O}\subset su(n)^{*}$ produces the symplectic leaf of the Poisson space $T^{*}SU(n)/SU(n)$ which is
+(8) $\mathcal{S}(\mathcal{O})=\mu^{-1}(\mathcal{O})/G=\{(x,g)\in su(n)\times SU(n)\mid x-g^{-1}xg\in\mathcal{O}\}/G.$
+Here $\mu:T^{*}G\to\mathfrak{g}^{*}$ is the moment map for the adjoint action, $\mu(x,g)=x-g^{-1}xg$.
+When $\mathcal{O}$ is a coadjoint orbit of rank one, the corresponding Hamiltonian reduction is isomorphic to $T^{*}H_{\mathrm{reg}}/S_{n}$ as a symplectic manifold.
+Here $H_{\mathrm{reg}}=\{\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n});\ \gamma_{i}\neq\gamma_{j}\}$ [^3] Clearly $T^{*}H_{\mathrm{reg}}/S_{n}$ is a smooth symplectic manifold..
+A rank $1$ orbit in $su(n)$ consists of traceless Hermitian matrices of rank $1$. They can be parametrized as
+(9) $\mu_{ij}=\overline{\varphi_{i}}\varphi_{j}-c\delta_{ij},\quad c=\frac{1}{n}\sum_{i=1}^{n}|\varphi_{i}|^{2}$
 modulo transformations $\varphi_{j}\mapsto e^{i\theta}\varphi_{j}$.
-It is clear that the second Casimir function 111 The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (5) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
-$\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
-The moment map brings the $SU(n)$-orbit through $(x,g)\in T^{*}G$ to
+It is clear that the second Casimir function  [^4] The values of higher Casimir functions are the rank $1$ orbit are determinated by $c$ and (10) $\mathrm{Tr}(\mu^{k})=c^{k}\,(n-1)\left((n-1)^{k-1}+(-1)^{k}\right).$ is
+(11) $\mathrm{tr}(\mu^{2})=c^{2}n(n-1).$
+The moment map for the adjoint action of $G$ on brings $(x,g)\in T^{*}G$ to
 (12) $\mu=x-g^{-1}xg.$
-From the construction of the phase space as a Hamiltonian reduction of $T^{*}SU(n)$ with respect to rank $1$ orbit the functions
-(7) $H_{k}=\frac{1}{k}\mathrm{Tr}\big{(}x^{k}\big{)},\ \ k=2,\dots,n$
-on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$ Poisson commute.
-This follows immediately from the symplectic structure on $T^{*}G=\{(x,g)\in\mathfrak{su}(n)\times SU(n)$.
-As a consequence the Poisson commute on $\mathcal{S}(\mathcal{O})$.
+Functions
+(13) $H_{k}(x,g)=\frac{1}{k}\mathrm{Tr}\big(x^{k}\big),\ \ k=2,\dots,n$
+on $T^{*}G$ Poisson commute.
+This follows immediately from the symplectic structure on $T^{*}G$.
+Because they are $G$-invariant, they are functions on $T^{*}G/G$.
+As a consequence, their restriction to $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G\subset T^{*}G/G$ also Poisson commute.
 Functions $\{H_{k}\}_{k=2}^{n}$ are independent on $S(\mathcal{O})$ and because $\mathrm{dim}\,S(\mathcal{O})=2n-2$ they define a Liouville integrable system.
 2.2. The diagonalization of $g$ and natural coordinates
-Let us briefly recall the Hamiltonian reduction construction of Calogero-Moser system by Kazhdan, Kostant, and Sternberg [4], extended to Calogero-Moser-Sutherland.
-We can choose a representative of the $SU(n)$–orbit through $(x,g)$ in (4) such that $g$ is diagonal:
+Let us recall how to obtain the Hamiltonians of the CMS system.
+Choose a representative of the SU(n)SU(n)–orbit through $(x,g)$ in (8) such that $g$ is diagonal:
 (14) $g=\mathrm{diag}(\gamma_{1},\ldots,\gamma_{n}).$
@@
 Thus $|\varphi_{i}|^{2}=c,\ c>0$.
-Note that if $\gamma_{i}=\gamma_{j}$, the equation (9) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus $\gamma_{i}\neq\gamma_{j}$.
-When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (8) is the Cartan subgroup of $SU(n)$ of unitary diagonal unimodular matrices.
+Note that if $\gamma_{i}=\gamma_{j}$, the equation (15) implies $\overline{\varphi}_{i}\varphi_{j}=0$ but this contradicts to $\lvert\varphi_{i}\rvert^{2}=c$. Thus, $\gamma_{i}\neq\gamma_{j}$.
+When $\gamma_{i}\neq\gamma_{j}$ the stabilizer of (14) is the Cartan subgroup of $SU(n)$, i.e.
+the subgroup of unitary diagonal unimodular matrices.
 By the action of this subgroup on $x$ we can bring $\varphi_{j}$ to the form $\varphi_{j}=\sqrt{c}\,e^{i\theta}$.
-This gives the description of the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $(x_{ii},\gamma_{i})$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
+This describes the phase space as the quotient of $T^{*}H_{reg}$ (with coordinates $x_{ii}$ on $\mathfrak{g}^{*}$, and $\gamma_{i}$ on $H_{reg}\subset G$) with respect to the action of $S_{n}$ by permutation of pairs $(x_{ii},\gamma_{i})$.
 For $j\neq k$ we have
 (17) $x_{jk}=\frac{c}{1-\gamma_{j}^{-1}\gamma_{k}}.$
-As it should be, we have $\bar{x}_{kj}=x_{jk}$.
 Setting $p_{i}:=x_{ii}$ we have
 (18) $x=\begin{pmatrix}p_{1}&\frac{c}{1-\gamma_{1}^{-1}\gamma_{2}}&\ldots&\frac{c}{1-\gamma_{1}^{-1}\gamma_{n}}\\ \frac{c}{1-\gamma_{2}^{-1}\gamma_{1}}&p_{2}&\ldots&\ldots\\ \ldots&\ldots&\ldots&\ldots\\ \frac{c}{1-\gamma_{n}^{-1}\gamma_{1}}&\ldots&\ldots&p_{n}\end{pmatrix}$.
-The unimodularity implies $\sum_{i=1}^{n}p_{i}=0$. For the first two Hamiltonians we have:
+The tracelessness of $x$ implies $\sum_{i=1}^{n}p_{i}=0$, and the unimodularity of $\gamma$ implies $\gamma_{1}\dots\gamma_{n}=1$. For the first two Hamiltonians, we have:
 (19) $H_{2}=\frac{1}{2}\mathrm{Tr}\big(x^{2}\big)=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{2}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{n}\frac{c^{2}}{(1-\gamma_{j}^{-1}\gamma_{i})(1-\gamma_{i}^{-1}\gamma_{j})}=\frac{1}{2}\sum_{i=1}^{n}p_{i}^{2}+\frac{1}{4}\sum_{\begin{subarray}{c}i<j\end{subarray}}^{n}\frac{c^{2}}{\sin^{2}(\tfrac{q_{i}-q_{j}}{2})},$
@@
 Here $\gamma_{j}=\exp(iq_{j})$.
-2.3. The diagonalization of $x$ and the action-angle variables
-Geometrically, any integrable system is a projection from a $2n$-dimensional symplectic manifold to an $n$-dimensional space, such that the generic fiber is Lagrangian.
-For Calogero–Moser systems we have the projection
-$\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(x,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
-Here $x$ is given by (12) as the function of $p$ and $\gamma$, and for $\mathfrak{su}(n)$.
-$W=S_{n}$, and $[x]$ is a collection of eigenvalues of $x$.
-The quotient space $\mathfrak{h}^{*}/W$ can be modeled as:
-$\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}.$
-The main observation of this paper is that the image of this mapping is a shifted cone in $\mathbb{R}^{n-1}$.
-Preimages of strata of this cone stratify $T^{*}H_{\mathrm{reg}}$.
-For CMS model we describe the strata later in this section.
-2.3.1. The generic stratum
-Any $x\in\mathfrak{su}(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$. let us choose a representative $(x,g)$ in our phase space where $x$ is diagonal:
-(15) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset\mathfrak{su}(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
-The Hamiltonians, evaluated in terms of eigenvalues of $x$ are simply symmetric power functions:
-$H_{k}=\sum_{j=1}^{n}x_{j}^{k}.$
+The Hamiltonian reduction and the Poisson structure on $T^{*}G$ imply that $G$-invariant functions on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ that come from $G$-invariant functions on $\mathfrak{g}^{*}$ form a Poisson commutative subalgebra.
+2.3. The CMS system as a Lagrangian fibration
+Geometrically, any integrable system is a projection from a 2n2n–dimensional symplectic manifold to an nn–dimensional space, such that the generic fiber is Lagrangian.
+For the CMS system this projection is
+(21) $\pi:\mathcal{S}(\mathcal{O})\to{su(n)^{*}}/SU(n)\simeq\mathfrak{h}^{*}/W.$
+Here $\mathcal{S}(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ is the phase space obtained by the Hamiltonian reduction, and $\mathfrak{h}\subset{su(n)}$ is the Cartan subgroup (real traceless diagonal matrices), and the Weyl group $W$ is the symmetric group $S_{n}$.
+From the results of the previous section, the CMS Lagrangian projection is equivalent to
+(22) $\pi\colon T^{*}H_{\mathrm{reg}}\to\mathfrak{h}^{*}/W,\quad(p,\gamma)\mapsto[x]\in\mathfrak{g}^{*}/G\simeq\mathfrak{h}^{*}/W.$
+Here $x$, as the function of $p$ and $\gamma$, is given by (18). For $su(n)$ we can identify $[x]$ with an ordered collection $(x_{1}\geq x_{2}\geq\dots\geq x_{n})$ of eigenvalues of $x$:
+(23) $\mathfrak{h}^{*}/W\simeq\left\{x\in\mathbb{R}^{n}\ \middle|\ x_{1}+\dots+x_{n}=0,\ x_{1}\geq\dots\geq x_{n}\right\}\subset\mathbb{R}^{n-1}.$
+This is the principal Weyl chamber for $su(n)$.
+Now let us describe the image of this projection as a shifted principal Weyl chamber in $\mathbb{R}^{n-1}$ and the stratification of $T^{*}H_{\mathrm{reg}}$ by the primages of walls of the shifted principal Weyl chamber.
+3. The base of the Lagrangian projection for the CMS model and the stratification of $S(\mathcal{O})$
+3.1. Solutions to the moment map equation
+Here we will describe the image $B(\mathcal{O})\subset\mathfrak{h}^{*}/W$ of the Lagrangian projection (22).
+We show that it is a convex set which can be regarded as a stratified space, which is a shifted principal Weyl chamber.
+We also describe the induced stratification of the phase space $S(\mathcal{O})$.
+Any $x\in su(n)^{*}$ is diagonalizable by the conjugation action of $SU(n)$.
+Let us choose a representative $(x,g)\in\mathfrak{g}^{*}\times G$ of the point $G(x,g)\in\mu^{-1}(\mathcal{O})/G\subset T^{*}G/G$ in the phase space where $x$ is diagonal:
+(24) $x=\mathrm{diag}\left(x_{1},\ldots,x_{n}\right)\in\mathfrak{h}^{*}\subset su(n)^{*},\qquad x_{1}+x_{2}+\ldots+x_{n}=0,\qquad x_{1}\geq x_{2}\geq\dots\geq x_{n}.$
 Theorem 1.
-Solutions to the momentum map equation
-(16) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\cdot 1.$
-can be described in the following way. Let’s $\lambda=g\bar{\psi}$, then the following holds
-1) The following conditions are equivalen
+Solutions to the momentum map equation [^5] From now on, when it is not ambiguous, we will write $a$ instead of $a\cdot\text{Id}$ for the matrix which is a multiple of the identity matrix.
+(25) $x-g^{-1}xg=\bar{\psi}\psi^{t}-c\ \$
+can be described in the following way. Define $\lambda=g\bar{\psi}$, then the following holds:
+1) The following conditions are equivalent
   • $x_{i}-x_{i+1}\geq c$ for all $i=1,\ldots,n-1$,
@@
 4) For $x_{i}-x_{i+1}>c$ with $i=1,\ldots,n-1$ matrix $g$ has form
-(18) $g_{ij}=\frac{\lambda_{i}\,\psi_{j}}{x_{j}-x_{i}+c},$
+(27) $g_{ij}=\frac{\lambda_{i}\,\psi_{j}}{x_{j}-x_{i}+c}\ \ .$
 Proof.
-1) The equation for the moment map can be written as
-(19) $g^{-1}xg=x+c\cdot 1-\bar{\psi}\psi^{t},$
-Following lemma 3.4 from [3], let’s consider the equality of characteristic polynomials of both parts of the equation (31)
-(20) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c\cdot 1-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
-Here we take into account that for diagonal matrix $A$
-(21) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{adj}u,$
-where $A^{adj}$ is the adjugate matrix to $A$, which defined by $A^{adj}A=\mathrm{det}A\cdot 1$. This is a well-known fact which follows from Lemma 1 in the appendix A.
-Equation (20) at the point $z=x_{i}+c$ is
+Let us start with part 1). The equation for the moment map can be written as
+(28) $g^{-1}xg=x+c-\overline{\psi}\psi^{t}.$
+Let us consider the equality of characteristic polynomials of both parts of this equation [^6] Compare with Lemma 3.4 from [3].,
+(29) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x+c-\bar{\psi}\psi^{t}-z)=\prod_{j=1}^{n}(x_{j}+c-z)-\sum_{j=1}^{n}|\psi_{j}|^{2}\prod_{k\neq j}(x_{k}+c-z).$
+Here we take into account that for a diagonal matrix AA
+(30) $\mathrm{det}\left(A+uv^{t}\right)=\mathrm{det}\left(A\right)+v^{t}A^{\text{adj}}u,$
+where $A^{\text{adj}}$ is the adjugate matrix to $A$, which defined by $A^{\text{adj}}A=\mathrm{det}A\cdot{\rm Id}$. This is a well-known fact which follows from lemma 1 in appendix A.
+The identity (29) for $z=x_{i}+c$ is
 (31) $\prod_{j=1}^{n}(x_{i}-x_{j}+c)=|\psi_{i}|^{2}\prod_{j\neq i}(x_{i}-x_{j}).$
-It follows from the equation (22) for $i=1$ that $|\psi_{1}|$ is strictly positive
+which proves (26).
+The equation (31) for $i=1$ implies that $|\psi_{1}|$ is strictly positive
 (32) $|\psi_{1}|^{2}=c\prod_{j>1}\frac{x_{1}-x_{j}+c}{x_{1}-x_{j}}>0.$
-In equation (22) for $i=2$
-(24) $c\ \frac{(x_{2}-x_{1}+c)}{x_{2}-x_{1}}\prod_{j>2}\frac{(x_{2}-x_{j}+c)}{(x_{2}-x_{j})}=|\psi_{2}|^{2}$
-the product over $j>2$ is strictly positive. So we have
-(25) $|\psi_{2}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{1}\geq x_{2}+c,$
-where equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
-Continuing this reasoning, we get the statement of the proposition. So, we have
-(26) $|\psi_{k}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
-where equality $x_{k-1}=x_{k}+c$ is achieved if $|\psi_{k}|^{2}=0$.
+In the equation (31) for $i=2$ the product over $j>2$ is strictly positive:
+[…]
+This implies that
+(34) $|\psi_{2}|^{2}\geq 0$
+if and only if $x_{1}\geq x_{2}+c$ and the equality $x_{1}=x_{2}+c$ is achieved if and only if $|\psi_{2}|^{2}=0$.
+Continuing this reasoning, we get:
+(35) $|\psi_{k}|^{2}\geq 0\ \$
+if and only if $x_{k-1}\geq x_{k}+c$ and the equality $x_{k-1}=x_{k}+c$ is achieved only if $|\psi_{k}|^{2}=0$.
 The equation for the moment map can be rewritten for $\lambda$ as
-(27) $gxg^{-1}=x-c\cdot 1+\lambda\bar{\lambda}^{t},$
-Equality of characteristic polynomials gives
-(28) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c\cdot 1+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
-Equation (28) at the point $z=x_{i}-c$ is
+(36) $gxg^{-1}=x-c+\lambda\bar{\lambda}^{t}.$
+The equality of characteristic polynomials gives
+(37) $\prod_{j=1}^{n}(x_{j}-z)=\mathrm{det}(x-c+\lambda\bar{\lambda}^{t}-z)=\prod_{j=1}^{n}(x_{j}-c-z)+\sum_{j=1}^{n}|\lambda_{j}|^{2}\prod_{k\neq j}(x_{k}-c-z).$
+Evaluating this idenity (37) at the point $z=x_{i}-c$ gives
 (38) $\prod_{j=1}^{n}(x_{j}-x_{i}+c)=|\lambda_{i}|^{2}\prod_{j\neq i}(x_{j}-x_{i}).$
 Using a similar argument as for equation (31), we get $|\lambda_{n}|^{2}>0$ and
-(30) $|\lambda_{k-1}|^{2}\geq 0\qquad\Leftrightarrow\qquad x_{k-1}\geq x_{k}+c,$
-where equality $x_{k-1}=x_{k}+c$ is achieved if $|\lambda_{k-1}|^{2}=0$.
-2) This part of the statement directly follows from (22) and (29).
-3) After the multiplication of the momentum map (31) by $g$ the equation for the moment map becomes
-(31) $gx+cg-xg=\lambda\psi^{t},$
-$(i,j)$ whose element has the form
+(39) $|\lambda_{k-1}|^{2}\geq 0$
+if and only if $x_{k-1}\geq x_{k}+c$, where equality $x_{k-1}=x_{k}+c$ is achieved only if $|\lambda_{k-1}|^{2}=0$. This proves part 1).
+Part 2) directly follows from (31) and (38). We already proved 3).
+Let us prove 4). After the multiplication of the momentum map (40) by $g$, the equation for the moment map becomes
+(40) $gx+cg-xg=\lambda\psi^{t}\ \ .$
+Matrix elements of this equation are a system of equations for $g_{ij}$:
 (41) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
-which fix the form of the thirt part of the statement
-which implies (18).
-Remark 1.
+which implies (27). ∎
+Corollary 1.
+The image of the projection (22) is the subset $B(\mathcal{O})\subset\mathbb{R}^{n-1}/S_{n}$ which can be described as
+(42) $B(\mathcal{O})=\{(x_{1},\dots,x_{n})\mid x_{i}-x_{i+1}\geq c,\ \sum_{i=1}^{n}x_{i}=0\}.$
+This space is naturally stratified. Let $\{j\}=\{j_{0},j_{1},\dots,j_{s}\}$ be a collection where $j_{0}=1,1<j_{a}<j_{a+1}\leq n-1$. Then
+(43) $B(\mathcal{O})=\sqcup_{\{j\}}B(\mathcal{O})_{\{j\}},$
+where
+[…]
+It is easy to check that
+(45) $\overline{B(\mathcal{O})_{\{j\}}}=\sqcup_{\{k\}\leq\{j\}}B(\mathcal{O})_{\{k\}}.$
+Here $\{k\}\leq\{j\}$ means that $\{k\}$ is a subset of $\{j\}$. This stratification for $G=SU(3)$ is illustrated on Figure 1.
+This stratification of $B(\mathcal{O})$ induces a stratification of $S(\mathcal{O})$ by the primages of $B$-strata, which we will describe below.
+3.2. The maximal stratum
 If $x_{i}>x_{i+1}+c$, the stabilizer of $x$ is the Cartan subgroup in $SU(n)$, i.e., unitary diagonal matrices.
@@
 (46) $g=\alpha X,\qquad X_{ij}=\frac{|\lambda_{i}||\psi_{j}|}{x_{j}-x_{i}+c},\qquad\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,$
-where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. with $\mathrm{det}\,X=1$ according to
-The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
+where $X$ is orthogonal matrix, i.e. $X^{t}=X^{-1}$. The Cauchy determinant formula implies that $\mathrm{det}\,X=1$, and therefore $\alpha_{1}\cdot\ldots\cdot\alpha_{n}=1$.
 Remark 1.
 We can multiply $g$ by a central element $\omega\in Z(SU(n))$:
-(34) $\omega=\exp\left(\frac{2\pi i}{n}k\right)\cdot 1_{n},\qquad\text{ where }k=0,\ldots,n-1.$
+(47) $\omega=\exp\left(\frac{2\pi i}{n}k\right){\rm Id},\qquad\text{ where }k=0,\ldots,n-1.$
 If $\gamma_{j}$ are eigenvalues of the matrix $g$, such multiplication changes $\gamma_{j}\mapsto\omega\gamma_{j}$.
-Eigenvalues of $g$ are exponents of coordinates of Calogero–Moser–Sutherland model $\gamma_{j}=\exp(iq_{j})$.
+Eigenvalues of $g$ are exponents of coordinates of the CMS model $\gamma_{j}=\exp(iq_{j})$.
 Thus the multiplication of $g$ by $\omega$ shifts coordinates as
+(48) $q_{j}\mapsto q_{j}+\frac{2\pi}{n}k\ (\mathrm{mod}\ 2\pi)\ .$
+Thus, we described the maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}\subset S(\mathcal{O})$.
+It has a natural coordinate system, $\{x_{i}\}$ that lie inside of the shifted principal Weyl chamber $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\{\alpha_{i}\}$, $|\alpha_{i}|=1$ and $\prod_{i=1}^{n}\alpha_{i}=1$.
+3.3. Low–dimensional strata $S(\mathcal{O})_{\{j\}}$
+To describe all strata, we need the following technical fact.
+Proposition 1.
+If $\psi_{k}=0$ for some $k=2,\dots,n$, then the only nonzero matrix element of $g$ in the (k1)(k-1)-th row and the $k$-th column is $g_{k-1,k}$ and $|g_{k-1,k}|=1$.
+Proof.
+The equation (41) for $j=k$ becomes
+(49) $(x_{k}-x_{i}+c)g_{ik}=0.$
+If $g_{ik}=0$ for all $i$, then $g$ not invertible.
+Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$.
+Thus, in the equation above $g_{k-1,k}$ can be non-zero.
+For $j\neq k-1$ we have $x_{j}-x_{k}+c\neq 0$, and therefore $g_{jk}=0$.
+This means that the $k$-th column has only one non-zero element $g_{k-1,k}$.
+The unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the (k1)(k-1)-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$.
+∎
+Fix numbers $\{j\}=\{j_{0}=1,j_{1},\dots,j_{s}\}$ such that $1<j_{1}<j_{2}<\dots<j_{s}\leq n$.
+Definition 1.
+The stratum $S(\mathcal{O})_{\{j\}}\subset S(\mathcal{O})$ is the subspace where $\psi_{k}\neq 0$ when $k\in\{j_{0},j_{1},\dots,j_{s}\}$ and $\psi_{k}=0$ otherwise.
+Equivalently, $x_{k}-x_{k+1}>c$ if $k\in\{j_{0},j_{1},\dots,j_{s}\}$ and $x_{k}-x_{k+1}=c$ otherwise.
+It is clear that $S(\mathcal{O})_{\{j\}}=\pi^{-1}(B(\mathcal{O})_{\{j\}})$.
+Define $\widehat{g}=P^{-1}g$, where $P$ is the signed matrix of cyclic permutation
+(50) $P=\begin{pmatrix}0&-1&0&\ldots&0\\ 0&0&-1&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&-1\\ 1&0&0&\ldots&0\\ \end{pmatrix}$.
+Proposition 1 implies
+(51) $\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
+if neither $i$ nor $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
+The moment map equation (12) give the following identity
+(52) $x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c,$
+where $x^{P}=P^{-1}xP$ is the diagonal matrix with entries $x_{j}^{P}=x_{j-1}$ and $x_{0}=x_{n}$. For matrix elements of $\widehat{g}$ we obtain equations
+(53) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}.$
+Define
+(54) $\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
+Theorem 2.
+The following holds:
+  • $\widehat{\lambda}_{i}=0$, if $i\neq j_{a}$, $a=0,\dots,s$.
+  • (55) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})}.$
+  • (56) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}.$
+  • By the action of the Cartan part of the stabilizer of $x$, we can bring $\widehat{g}$ to the form
+    (57) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}.$
+Proof.
+Write the moment map equation as
+(58) $\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
+For the characteristic polynomial of the right side, using lemma 1 from appendix A we obtain
+(59) $\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
+Evaluation of this identity at $z=x_{j_{a}}+c$ gives
+(60) $\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
+which proves the formula (55) for $|\psi_{j_{a}}|^{2}$.
+Similarly, writing the moment map equation as
+(61) $\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
+and using lemma 1 gives the proof of the formula (56) for $|\widehat{\lambda}_{j_{a}}|^{2}$.
+Finally, the moment map equations (53) give
+(62) $\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
+By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives formula (57). ∎
+Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer of $x$.
+Corollary 2.
+We can choose a representative $(x,g)$ of the conjugacy class $(x,g)]\in\mu^{-1}(\mathcal{O})/G$, where $x$ is diagonal as above and $g$ is the product
+(63) $g=\alpha\,P\,Y.$
+Here $P$ is the cyclic permutation matrix (50), $Y$ is
+$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\},$
+(64) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad\,\ i\notin\{j_{0},\ldots,j_{s}\},$
+(65) $Y_{j_{a}j_{b}}$ $=(-1)^{1-\delta_{a,0}}\frac{|\widehat{\lambda}_{j_{a}}||\psi_{j_{b}}|}{x_{j_{b}}-x_{j_{a}}^{P}+c},$
+and $\alpha$ is the following diagonal matrix
+(66) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a}.$
+Indeed, let $\beta$ be an element of the stabilizer of $x$, i.e., a diagonal unitary matrix. We already used transformations
+(67) $g\mapsto\beta g\beta^{-1}$
+with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real. Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
+(68) $\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
+We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$. This proves corollary 2.
+One can show that
+(69) $\mathrm{det}\,Y=1$
+and that the real-valued matrix $Y$ is orthogonal
+(70) $YY^{t}={\rm Id}.$
+Note that because $\det(P)=1$, we have $\det(\alpha)=1$.
+Corollary 3.
+$S(\mathcal{O})_{\{j\}}\simeq\mathbb{T}^{s}\times\mathbb{R}^{s}$ with $x_{j_{a}}$ being coordinates on $\mathbb{R}^{s}$ and $\alpha_{j_{a}}$ being coordinates on $\mathbb{T}^{s}$.
+3.4. The dynamical stratification of the phase space
+The results of the previous sections imply that we have a stratification of $S(\mathcal{O})$ given by the preimages of walls of the shifted Weyl chamber:
 […]
-Thus, we described coordinates on the generic (maximal) stratum: real coordinates $\{x_{i}\}$ and $\{\mathrm{arg}\,\alpha_{i}\}$ with satisfying constrains $x_{i}>x_{i+1}+c,\ \sum_{i=1}^{n}x_{i}=0$ and $\prod_{i=1}^{n}\alpha_{i}=1$, $|\alpha_{i}|=1$.
-Now let us describe all strata and let us show that $(x,\alpha)$ are the action-angle variables.
-Define $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
+The first part in this disjoint union is the fixed point $S(\mathcal{O})_{\{1\}}$ corresponding to $s=0$.
+Each stratum $S(\mathcal{O})_{\{j_{0},\dots,j_{s}\}}$ has dimension $2s$.
+The maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ has maximal dimension 2n22n-2.
+This is a stratification where strata are parametrized by the partially ordered set $A=\{1\}\cup\dots\cup\{1<j_{1}<\dots<j_{s}\}$ with the partial order $\{k\}\leq\{j\}$ if $\{k\}$ is a subset of $\{j\}$.
+, see corollary 3, and
+(72) $\overline{S(\mathcal{O})_{\{j\}}}=\sqcup_{\{k\}\leq\{j\}}S(\mathcal{O})_{\{k\}}.$
+This property can be checked by taking the appropriate limit. For example, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$
+(73) $|\psi_{i}|^{2}$ $\to 0,\qquad\qquad\text{for }i\neq j_{a},$
+(74) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{ for }i=j_{a}.$
+Here on the left hand side $|\psi_{i}|^{2}$ is taken from (26), on the right hand side $|\psi_{j_{a}}|^{2}$ is taken from (55), and
+(75) $|\lambda_{i}|^{2}$ $\to 0,\qquad\qquad\text{for }i\neq j_{a}-1,$
+(76) $|\lambda_{i}|^{2}$ $\to|\widehat{\lambda}_{j_{a}}|^{2},\qquad\text{ for }i=j_{a}-1\ \mathrm{mod}\ n.$
+3.4.1. Natural coordinates on $S(\mathcal{O})_{\{j\}}$
+On the the maximum stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ define variable $y_{i}=x_{i-1}-x_{i}$ for $j=2,\ldots,n$, such that $y_{i}\geq c$. Then
 (77) $x_{i}=\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}.$
-Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\log\alpha_{k}$ for $j=2,\ldots,n$. Then,
+Define also $\theta_{j}=\tfrac{1}{i}\sum_{k=1}^{j-1}\ln\alpha_{k}$ for $j=2,\ldots,n$. Then,
 (78) $\alpha_{k}=e^{i(\theta_{k+1}-\theta_{k})},\qquad k=1,\ldots n.$
 Here we assume that $\theta_{1}=\theta_{n+1}=0$.
-Thus, we get coordinates on the generic (maximal) stratum: real coordinates $(y_{2}\ldots,y_{n})$ satisfying constrains $y_{i}>c$ and $(\theta_{2}\ldots,\theta_{n})$.
-In section 4 we show that $(y,\theta)$ are the action-angle variables.
-3. Low–dimensional tori
-The moment map equation (6) when $x$ is diagonal becomes
-(38) $(x_{j}-x_{i}+c)g_{ij}=\lambda_{i}\psi_{j},$
-where $\lambda=g\bar{\psi}$.
-Proposition 1.
-If for some $k$ we have $\psi_{k}=0$, then in the $k$-th column and $(k-1)$-th row of the martrix $g$ only nonzero element is $g_{k-1,k}\neq 0$ and $|g_{k-1,k}|=1$.
-Proof.
-Consider equation (38) for $j=k$
-(39) $(x_{k}-x_{i}+c)g_{ik}=0.$
-If for all $i$ elements $g_{ik}=0$ then $g$ not invertible.
-Theorem 1 implies that $\psi_{k}=0$ if and only if $x_{k}-x_{k-1}+c=0$.
-For $j\neq k-1$ $x_{j}-x_{k}+c\neq 0$ and therefore $g_{jk}=0$.
-This means that $k$-th column has only one non-zero element $g_{k-1,k}$.
-Unitarity condition $\left(g^{{\dagger}}g\right)_{ik}=\sum_{j=1}^{n}\bar{g}_{ji}g_{jk}=\bar{g}_{k-1,i}g_{k-1,k}=\delta_{ik}$ implies that in the $(k-1)$-th row there is only one non zero element $g_{k-1,k}$, and $|g_{k-1,k}|=1$.
-∎
-Assume that $\psi_{j}\neq 0$ only for $j=1,j_{1},\dots,j_{s}$ where $j_{1}<j_{2}<\dots<j_{s}$. For convenience, define $j_{0}=1$.
-Define the matrix $\widehat{g}$ as $P\widehat{g}=g$, where $P$ is the matrix of cyclic permutation with sign
-(40) $P=\begin{pmatrix}0&-1&0&\ldots&0\\ 0&0&-1&\ldots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\ldots&-1\\ 1&0&0&\ldots&0\\ \end{pmatrix}.$
-Proposition 1 implies
-$\widehat{g}_{ij}=\delta_{ij}\alpha_{i},\quad|\alpha_{i}|=1$
-if neither $i$ no $j$ are not in $\{j_{a}\}_{a=0}^{s}$.
-The moment map equations give the following identity
-$x-\widehat{g}^{-1}x^{P}\widehat{g}=\overline{\psi}\psi^{t}-c$
-where $x_{j}^{P}=x_{j-1}$ with $x_{0}=x_{n}$, and equations
-(41) $\left(x_{j_{b}}-x^{P}_{j_{a}}+c\right)\widehat{g}_{j_{a}j_{b}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}\,\psi_{j_{b}}$
-Define
-$\widehat{\lambda}_{j_{a}}=\sum_{d=0}^{s}\widehat{g}_{j_{a}j_{d}}\,\overline{\psi}_{j_{d}}.$
-Theorem 2.
-The following holds:
-  • $\widehat{\lambda}_{i}\neq 0,\lambda_{i}$, for $i\neq j_{a}$, $a=0,\dots,s$.
-  • (42) $|\psi_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{a}}-x^{P}_{j_{d}}+c)}{\prod^{s}_{d\neq a}(x_{j_{a}}-x_{j_{d}})},$
-  • (43) $|\widehat{\lambda}_{j_{a}}|^{2}=\frac{\prod_{d=0}^{s}(x_{j_{d}}-x^{P}_{j_{a}}+c)}{\prod^{s}_{d\neq a}(x^{P}_{j_{d}}-x^{P}_{j_{a}})}$
-  • By the action of the Cartan part of the stabilizer of $x$, we can bring $\widehat{g}$ to the form
-    (44) $\widehat{g}_{j_{a}j_{b}}=\alpha_{j_{a+1}}\frac{|\widehat{\lambda}_{j_{a}}|\,|\psi_{j_{b}}|}{x_{j_{b}}-x^{P}_{j_{a}}+c}$
-Proof.
-Write the moment map equation as
-$\hat{g}^{-1}x^{P}\hat{g}=x-\overline{\psi}\psi^{t}+c.$
-For the characteristic polynomial of the right side, using lemma 1 from the appendix A we obtain
-$\prod_{d=0}^{s}(x_{j_{d}}^{P}-z)=\prod_{d=0}^{s}(x_{j_{d}}+c-z)-\sum_{b=0}^{s}|\psi_{j_{b}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq b\end{subarray}}^{s}(x_{j_{d}}+c-z).$
-Taking $z=x_{j_{a}}+c$ we obtain
-$\prod_{d=0}^{s}(x_{j_{d}}^{P}-x_{j_{a}}-c)=-|\psi_{j_{a}}|^{2}\prod_{\begin{subarray}{c}d=0\\ d\neq a\end{subarray}}^{s}(x_{j_{d}}-x_{j_{a}})$
-which proves the formula (42) for $|\psi_{j_{a}}|^{2}$.
-Similarly, writing the moment map equation as
-$\hat{g}\,x\,\hat{g}^{-1}=x^{P}+\widehat{\lambda}\,\widehat{\lambda}^{*}-c$
-and using lemma 1 gives the proof of the formula (43) for $|\widehat{\lambda}_{j_{a}}|^{2}$.
-Finally, the moment map equations (41) give
-$\widehat{g}_{j_{a}j_{b}}=\frac{\widehat{\lambda}_{j_{a}}\,\psi_{j_{b}}}{x_{j_{b}}-x_{j_{a}}^{P}+c}.$
-By the action of the stabilizer of $x$, which is the Cartan subgroup $SU(n)$, we can make $\psi_{j_{b}}$ real. This gives (44). ∎
-Note that to make $\psi_{j_{a}}$ real we need only a subgroup of the stabilizer.
-Remark 3.
-We can write $g$ as the product
-(45) $g=\alpha\,P\,Y,$
-where $P$ is the cyclic permutation matrix (40).
-Let $\beta$ be an element of the stabilizer $x$, i.e. a diagonal unitary matrix.
-We use transformations
-(46) $g\to\beta g\beta^{-1}$
-with $\beta_{i}=1$, when $i\neq j_{a}$, to make $\psi_{j_{a}}$ real.
-Let $\beta^{\prime}\in H$ be such that $\beta^{\prime}_{i}=1$, when $i=j_{a}$. We have
-$\beta^{\prime}g\beta^{\prime-1}=\beta^{\prime}\alpha PY\beta^{\prime-1}=\beta^{\prime}\alpha\beta^{\prime-1}_{P}PY.$
-We can use these transformations to bring $\alpha$ to the form $\alpha_{i}=1$, $i\neq j_{a}$.
-Thus, we can fix the section of the action of the stabilizer group $H\subset SU(n)$:
-(47) $g=\alpha\,P\,Y,$
-where
-$Y_{ij}$ $=\delta_{ij},\quad i,j\notin\{j_{0},\ldots,j_{s}\}$
-(48) $Y_{i,j_{a}}=Y_{j_{a},i}$ $=0,\quad i\notin\{j_{0},\ldots,j_{s}\}$
-(49) $Y_{j_{a}j_{b}}$ $=(-1)^{1-\delta_{a,0}}\frac{|\widehat{\lambda}_{j_{a}}||\psi_{j_{b}}|}{x_{j_{b}}-x_{j_{a}}^{P}+c},$
-(50) $\alpha_{ij}=\alpha_{i}\delta_{ij},\qquad|\alpha_{i}|=1,\qquad\text{and}\quad\alpha_{i}=1\ \text{ for }\ i\neq j_{a},$
-and $P$ is defined (40).
-Note that with this choice, we have
-(51) $\mathrm{det}\,\alpha=\mathrm{det}\,P=\mathrm{det}\,Y=1$
-and real matrix $Y$ is othogonal
-$YY^{t}=I.$
-Define $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1},$ then due to (36)
-(52) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c).$
-Thus, we described coordinates $(y_{j_{a}},\alpha_{j_{a}})$ on $s$-dimention strata: $y_{j_{a}}>c$ and $\{\alpha_{j_{a}}\}_{a=0}^{s}$ with constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$.
+This gives a convenient set of coordinates on the maximal stratum: real coordinates $(y_{2},\ldots,y_{n})$ with constraints $y_{i}>c$ and coordinates $(\theta_{2},\ldots,\theta_{n})$ on unit tori.
+In section 4 we will show that $(y,\theta)$ are the action-angle variables.
+On $S(\mathcal{O})_{\{j\}}$ define variables $y_{j_{a}}=x_{j_{a}}-x_{j_{a}-1}$ then
+(79) $x_{i}=\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c),$
+where $\delta_{i\geq j}$ is a step function, $\delta_{i\geq j}=0$ if $i<j$ and $\delta_{i\geq j}=1$ if $i\geq j$.
+Set $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ and define $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$.
+Thus, we described coordinates $(y_{j_{a}},\theta_{j_{a}})$ with $y_{j_{a}}>c$ on $S(\mathcal{O})_{\{j\}}$.
+We will show that these are action-angle variables on this stratum.
+Because $\{y\}$ are coordinates on the base $B(\mathcal{O})_{\{j\}}$ of the Lagrangian fibration (22) of $S(\mathcal{O})_{\{j\}}$, they can be taken as action variables.
+In the next section we will show that $\{\theta\}$ are the corresponding angle variables.
+3.4.2. Two examples of low–dimensional strata
 Example 1.
-(The fixed point). In the case when all $\psi_{i}=0$ for $i\geq 2$, we have
+The fixed point $S(\mathcal{O})_{\{1\}}$. In this case all $\psi_{i}=0$ for $i\geq 2$, i.e., $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$, and we have
   • $|\psi_{1}|^{2}=|\hat{\lambda}_{1}|^{2}=nc$.
   • $x_{i}=\frac{n-1}{2}c-(i-1)c$ for $1=2,\ldots,n$.
-(53) $g=P.$
+  • $g=P.$
 In terms of original CMS coordinates we have $q_{j}=\frac{2\pi j}{n}$.
-The action of the center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}\mod(2\pi)$.
+The center of $SU(n)$, $g\mapsto\omega^{k}g$, $\omega^{n}=1$, acts by the cyclic shift $q_{j}\mapsto q_{j}+\frac{2\pi k}{n}({\rm mod}\ 2\pi)$.
 Example 2.
-(One-domensional tori). Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ when $i\neq 1,j$.
+Two–dimensional strata $S(\mathcal{O})_{\{1,j\}}$. Fix $2\leq j\leq n$ and assume that $\psi_{i}=0$ only when $i\neq 1,j$. The following is a consequence of theorem 2 and corollary 2
   • $|\psi_{1}|^{2}+|\psi_{j}|^{2}=|\hat{\lambda}_{1}|^{2}+|\hat{\lambda}_{j}|^{2}=nc$.
-  • For $y=x_{j-1}-x_{j}$, $y>c$; $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$, and with $\sum_{i=1}^{n}x_{i}=0$:
-    (54) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c),$
-    where $\delta_{i\geq j}$ is a step function.
-  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j},1,\ldots,1\right)$.
-(55) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
-$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c}$
+  • $x_{i-1}-x_{i}=c$ for $i=2,\ldots,n$ and $i\neq j$.
+  • Define $y=x_{j-1}-x_{j}$, then $y>c$. Together with $\sum_{i=1}^{n}x_{i}=0$ this gives:
+    (80) $x_{i}=\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y-c).$
+  • $g=\alpha PY$, where $\alpha=\mathrm{diag}\left(\alpha_{1},1,\ldots,1,\alpha_{j}=\alpha_{1}^{-1},1,\ldots,1\right)$.
+(81) $|\psi_{1}|^{2}$ $=(j-1)c\,\frac{y+(n-1)c}{y+(j-2)c},\qquad\ \quad|\psi_{j}|^{2}=(n-j+1)c\,\frac{y-c}{y+(j-2)c},$
+$|\hat{\lambda}_{1}|^{2}$ $=(n-j+1)c\frac{y+(n-1)c}{y+(n-j)c},\qquad|\hat{\lambda}_{j}|^{2}=(j-1)c\ \frac{y-c}{y+(n-j)c},$
 and
-(56) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix},$
+(82) $Y=\begin{pmatrix}r&0&\ldots&q&\ldots&0\\ 0&1&\ldots&0&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ -q&0&\ldots&r&\ldots&0\\ \vdots&\vdots&\ddots&\vdots&\ddots&\vdots\\ 0&0&\ldots&0&\ldots&1\\ \end{pmatrix}\$,
 where
 (83) $r^{2}=\frac{(j-1)c\,(n-j+1)c}{(y+(j-2)c)(y+(n-j)c)},\qquad q^{2}=\frac{(y-c)(y+(n-1)c)}{(y+(j-2)c)(y+(n-j)c)}.$
-Note that $r^{2}+q^{2}=1$. It is clear that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.
-Remark 4.
-Theorem 2 can be obtained from theorem 1 directly using the appropriate limit. In particular, if $x_{i-1}\to x_{i}+c$ for $i\neq j_{a}$, $a=0,\dots,s$ ($j_{0}=1$)
-(58) $|\psi_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a},$
-(59) $|\psi_{i}|^{2}$ $\to|\psi_{j_{a}}|^{2},\qquad\text{for }i=j_{a},$
-where in the LHS $|\psi_{i}|^{2}$ from (17) and in the RHS $|\psi_{j_{a}}|^{2}$ from (42), and
-(60) $|\lambda_{i}|^{2}$ $\to 0,\qquad\text{for }i\neq j_{a}-1,$
-(61) $|\lambda_{i}|^{2}$ $\to|\widehat{\lambda}_{j_{a}}|^{2},\qquad\text{for }i=j_{a}-1\ \mathrm{mod}\ n.$
+Note that $r^{2}+q^{2}=1$, which implies that $\mathrm{det}(\,Y)=1$, therefore $\mathrm{det}\,\alpha=\alpha_{1}\,\alpha_{j}=1$.
 4. Action-angle variables
-The cotangent bundle $T^{*}G$ carries the natural symplectic form
+Here we show that natural coordinates $(y,\theta)$ are Darboux action-angle variables on each $S(\mathcal{O})_{\{j\}}$.
+The cotangent bundle $T^{*}G$ carries the natural symplectic form, which, after the trivialization of $T^{*}G$ by right translations can be written in terms of $(x,g)\in su(n)^{*}\times SU(n)\simeq T^{*}SU(n)$ as
 (84) $\Omega=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right).$
-Here we used the trivialization $T^{*}G\simeq\mathfrak{g}^{*}\times G$ by right translation.
-The identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$ is given by the Killing form.
-We also assume that $G$ is a compact matrix group.
-This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. After the trivialization and identification $\mathfrak{g}^{*}\simeq\mathfrak{g}$, it acts as
-$h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
-The symplectic form (62) restricted to $\mu^{-1}(\mathcal{O})$ gives a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$.
+Here $x$ is an $n\times n$ Hermitian traceless matrix an $g$ is and unitary unimodular $n\times n$ matrix.
+This symplectic form is $G$-invariant with respect to the lift of the conjugation action to $T^{*}G$. It acts on $(x,g)$ as
+(85) $h\colon\quad x\mapsto hxh^{-1},\quad g\mapsto hgh^{-1}.$
+The restriction of the symplectic form (84) to $\mu^{-1}(\mathcal{O})$ defines a presymplectic form, which remains $G$-invariant, and becomes a symplectic form $\Omega_{\mathcal{O}}$ on $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$.
 We can model $S(\mathcal{O})$ as the space of pairs $(x,g)$, up to the action $H\subset SU(n)$, where $x=(x_{1}\geq x_{2}\geq\ldots\geq x_{n})$, $\sum_{i=1}^{n}x_{i}=0$, is diagonalized.
-On the big stratum $g=\alpha X$ (33), on others strata $g=\alpha PY$, see remark 3.
-Now, let us describe the restriction $\Omega_{\mathcal{O}}$ of the symplectic form (62) in terms of $(y,\theta)$ coordinates.
+On the maximal stratum $g=\alpha X$ (46), on other strata $g=\alpha PY$, see remark 2.
+Now, let us describe the symplectic form $\Omega_{\mathcal{O}}$ in terms of $(y,\theta)$ coordinates introduced in section 3.4.1.
 Proposition 2.
-On the generic stratum
-(63) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j}.$
+On the maximal stratum $S(\mathcal{O})_{\{1,2,\dots,n\}}$ the symplectic form $\Omega_{\mathcal{O}}$ is
+(86) $\Omega_{\mathcal{O}}=-\sum_{j=1}^{n}\mathrm{d}x_{j}\wedge\mathrm{d}\eta_{j},$
 where $x_{i}$ are eigenvalues of $x$ and $\alpha_{j}=\mathrm{exp}(i\eta_{j})$.
 Proof.
-Using the representation (33) for $\Omega$ form we get
+Evaluating the symplectic form $\Omega$ on the representative $(x,g=\alpha X)$ we obtain
 (87) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha X)\,X^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right).$
 The first term vanishes:
-(65) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0,$
-where we take into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$.
-Thus
+(88) $\mathrm{Tr}\left(x\,\mathrm{d}X\,X^{-1}\right)=\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}X_{ij}\cdot(X^{t})_{ji}=\frac{1}{2}\sum_{i,j=1}^{n}x_{i}\,\mathrm{d}(X_{ij})^{2}=\frac{1}{2}\sum_{i=1}^{n}x_{i}\,\mathrm{d}\sum_{j=1}^{n}(X_{ij})^{2}=0.$
+Here we took into account the orthogonality of $X$, which implies $\sum_{j=1}^{n}(X_{ij})^{2}=1$. Thus, on the maximal stratum
 (89) $\Omega_{\mathcal{O}}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}\alpha\,\alpha^{-1}\right)=i\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\alpha^{-1}_{i}\mathrm{d}\alpha_{i}=-\sum_{i=1}^{n}\mathrm{d}x_{i}\wedge\mathrm{d}\eta_{i}.$
-Taking into account constraints $\sum_{j=1}^{n}\mathrm{d}x_{j}=0$ and $\sum_{j=1}^{n}\mathrm{d}\eta_{j}=0$ we can write $\Omega_{\mathcal{O}}$ as
-(67) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j}.$
+After changing variables from $(x,\eta)$ to $(y,\theta)$ we can write $\Omega_{\mathcal{O}}$ as
+(90) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j},$
 where $y_{j}=x_{j-1}-x_{j}$ and $\theta_{j}=\sum_{i=1}^{j-1}\eta_{i}$. Thus, on the generic stratum $(y_{j},\theta_{j})$ are Darboux action-angle coordinates.
 Proposition 3.
-On a $2s$-dimensional stratum of $S(\mathcal{O})$ the symplectic form $\Omega_{\mathcal{O}}$ becomes
-(68) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}}.$
+On the 2s2s–dimensional stratum $S(\mathcal{O})_{\{j\}}$ we have
+(91) $\Omega_{\mathcal{O}}=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
 where $y_{j_{a}}=x_{j_{a}-1}-x_{j_{a}}$ and $\theta_{j_{a}}=\sum_{b=a}^{s}\eta_{j_{b}}$, with $\alpha_{j_{b}}=e^{i\eta_{j_{b}}}$.
 Proof.
-Using representation (47) for $\Omega$ form we get
+Writing $g=\alpha PY$ as in (63) for the restriction of $\Omega_{\mathcal{O}}$ we get
 (92) $\Omega_{\mathcal{O}}=\left.i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}g\,g^{-1}\right)\right|_{\mu^{-1}(\mathcal{O})}=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}(\alpha PY)\,Y^{-1}P^{-1}\alpha^{-1}\right)=i\,\mathrm{d}\mathrm{Tr}\left(x\,\mathrm{d}Y\,Y^{-1}\right)+i\,\mathrm{d}\mathrm{Tr}\left(x^{P}\,\mathrm{d}\alpha\,\alpha^{-1}\right),$
 where $x^{P}=P^{-1}xP$.
-Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrains $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies.
+Define $\alpha_{j_{a}}=e^{i\eta_{j_{a}}}$ for $a=0,\ldots,s$. Constrain $\prod_{a=0}^{s}\alpha_{j_{a}}=1$ implies
 (93) $\eta_{j_{0}}\equiv\eta_{1}=-\sum_{a=1}^{s}\eta_{j_{a}}.$
 Expression $\mathrm{Tr}\left(x^{P}\,\mathrm{d}Y\,Y^{-1}\right)=0$ for the same reasons as in equation (88) and $\Omega_{\mathcal{O}}$ becomes
-(71) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=\\ -\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=\\ -\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
+(94) $\Omega_{\mathcal{O}}=i\sum_{a=0}^{s}\mathrm{d}x^{P}_{j_{a}}\wedge\alpha_{j_{a}}^{-1}\mathrm{d}\alpha_{j_{a}}=$- $\sum_{a=1}^{s}\frac{1-j_{a}}{n}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\eta_{1}-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(\frac{n+1-j_{b}}{n}-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=$
+$-\sum_{a=1}^{s}\sum_{b=1}^{s}\left(1-\delta_{j_{b}-1\geq j_{a}}\right)\mathrm{d}y_{j_{b}}\wedge\mathrm{d}\eta_{j_{a}}=-\sum_{b=1}^{s}\mathrm{d}y_{j_{b}}\wedge\sum_{a=b}^{s}\mathrm{d}\eta_{j_{a}}.$
 Here we used (93). In term of $\theta_{j_{a}}=\sum_{a=b}^{s}\eta_{j_{b}}$ this gives (91). ∎
-Remark 5.
-Note that $\theta$’s defined in proposition 3 differ slightly from $\theta$’s defined in proposition 2. This difference arises due to the introduction of a shift matrix $P$ (40).
-The coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux action-angle coordinates on this $s$-dimensional stratum.
-5. The multitime dynamics of Calogero–Moser–Sutherland system
-5.1. The multi-time dynamics on $T^{*}G$.
-Let $v_{k}$ be the Hamiltonian vector field on $T^{*}G\simeq\mathfrak{g}^{*}\times G$ generated by $H_{k}=\frac{1}{k}\operatorname{tr}(x^{k})$, i.e.
-$i_{v_{k}}\Omega=-dH_{k}$, where $\Omega$ is the symplectic form (62) on $T^{*}G$.
-We have
-(72) $v_{k}=-i\sum_{i,j=1}^{n}\left(x^{k-1}g\right)_{ij}\frac{\partial}{\partial g_{ij}},$
-Because Hamiltonians $H_{k}$ Poisson commute on $T^{*}G$, vector fields $v_{k}$ also commute. Their multi-time flows are solutions to the system of differential equations:
-(73) $\frac{\mathrm{d}x({\bf t})}{\mathrm{d}t_{k}}$ $=0,$
-(74) $\frac{\mathrm{d}g({\bf t})}{\mathrm{d}t_{k}}$ $=ix({\bf t})^{k-1}g({\bf t}).$
-The initial value problem $x(0)=x,\,g(0)=g$ for these equations can be solved explicitly:
-(75) $x({\bf t})=x,\qquad g({\bf t})=e^{i\sum_{k=1}^{n}t_{k}\,x^{k-1}}g.$
-This dynamics preserves Hamiltonians $H_{k}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
-5.2. The multi-time dynamics on the generic stratum.
-In coordinates $(y,\theta)$ the symplectic form on $S(\mathcal{O})$ obtained by the Hamiltonian reduction from (62) is
-(76) $\Omega_{\mathcal{O}}=-\sum_{j=2}^{n}\mathrm{d}y_{j}\wedge\mathrm{d}\theta_{j},$
-which means Poisson brackets between coordinate function $\{y_{i}\}$ and $\{\theta_{i}\}$ are
-$\{y_{i},y_{j}\}=0,\qquad\{\theta_{i},\theta_{j}\}=0,\qquad\{\theta_{i},y_{j}\}=\delta_{ij}.$
-Any $G$-invariant function $H$ on $T^{*}G$ defines a $G$-invariant function on $\mu^{-1}(\mathcal{O})\subset T^{G}$ and a function $S(\mathcal{O})=\mu^{-1}(\mathcal{O})/G$ which we denote by same letter $H$.
-Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
-(77) $v_{k}=\sum_{i,j}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j}}\frac{\partial}{\partial\theta_{j}},\qquad\text{ for }k=2,\ldots,n.$
-Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
-(78) $v_{k}=\sum_{j=2}^{n}u^{k}_{j}\frac{\partial}{\partial\theta_{j}},$
+Thus, the coordinates $(y_{j_{a}},\theta_{j_{a}})$ are Darboux coordinates on $S(\mathcal{O})_{\{j\}}$.
+The coordinates $y_{j_{a}}$ are also coordinates on the stratum $B(\mathcal{O})_{\{j\}}$ of $B(\mathcal{O})$.
+Complementary coordinates $\theta_{j_{a}}$ are coordinates on the tori $\pi^{-1}(y)$ and are angle coordinates.
+Now let us show that the multi-time dynamics is linear in $\theta$.
+5. The multi-time dynamics of Calogero–Moser–Sutherland system
+5.1. The multi-time dynamics on $T^{*}G$
+Let $F$ be a polynomial $G$-invariant function on $\mathfrak{g}^{*}$ and $v_{F}$ be the Hamiltonian vector field generated by the pull-back of $F$ to $T^{*}G\simeq\mathfrak{g}\times G$ with respect to the projection to the first factor.
+For $G=SU(n)$
+(95) $v_{F}(x,g)=-i\sum_{k,j+1}^{n}(\nabla Fg)_{kj}\frac{\partial}{\partial g_{kj}}.$
+Here $\nabla F(x)$ is the gradient of $F$.
+It is an $su(n)$-valued function on $su(n)^{*}$, defined as $(y,\nabla F(x))=\frac{d}{dt}F(x+ty)|_{t=0}$, where $(y,x)={\rm tr}(xy)$.
+The flow line of this vector field, passing theorugh $(x,g)$ at $t=0$ is
+(96) $(x(t),g(t))=(x,e^{t\nabla F(x)}g)\ \ .$
+Any two such functions $F_{1},F_{2}$ Poisson commute with respect to the standard Poisson bracket on $T^{*}G$, $\{F_{1},F_{2}\}=0$.
+As a consequence, the corresponding vector field commute $v_{F_{1}},v_{F_{2}}]=0$ and thus, for any $k$ such functions $F_{1},\dots,F_{k}$ the multi-time flow is is defined with flow lines
+(97) $(x(\mathbf{t}),g(\mathbf{t}))=(x,e^{i\sum_{j=1}^{k}t_{j}\nabla F_{j}(x)}g)\ \ .$
+This dynamics preserves the pull-back to $T^{*}G$ of any $G$-invariant function $H$ on $\mathfrak{g}^{*}$ and the moment map $\mu(x({\bf t}),g({\bf t}))=\mu(x,g)=x-g^{-1}xg$.
+5.2. The multi-time dynamics on $S(\mathcal{O})_{\{j\}}$
+As we proved in Corollary 2 on the stratum $S(\mathcal{O})_{\{j\}}$, where $\{j\}=\{j_{0}=1<j_{1}<\dots<j_{s}\leq n\}$, the matrix $g$ can be written as
+(98) $g=\alpha PY.$
+In coordinates $(y,\theta)$, described earlier, the restriction (the pull-back) of the symplectic form $\Omega$ to this stratum is
+(99) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
+which means that Poisson brackets between coordinate functions are:
+(100) $\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
+For a function $F$ which is the pull-back to $T^{*}G$ of a $G$-invariant function on $\mathfrak{g}^{*}$, the Hamiltonian vector fields $v_{F}$ being restricted to $S(\mathcal{O})_{\{j\}}$ can be written as
+(101) $v_{F}^{\{j\}}=\sum_{i=1}^{n}\sum_{a=1}^{s}\frac{\partial F}{\partial x_{i}}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
+Using (77) we can express $x_{i}$ in terms of $y_{i}$. This gives
+(102) $v_{F}^{\{j\}}=\sum_{a=2}^{s}u^{F,\{j\}}_{a}(y)\frac{\partial}{\partial\theta_{j_{a}}},$
 where
-(79) $u^{k}_{j}=\sum_{i=1}^{n}\left(\frac{1}{n}\sum_{j=2}^{n}(n+1-j)y_{j}-\sum_{j=2}^{i}y_{j}\right)^{k-1}\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right).$
-Flow lines of these vector fields are linear in coordinates $\theta_{l}$. The multitime flow line with $y_{j}(0)=y_{j}$, $\theta_{j}(0)=\theta_{j}$ is
-(80) $y_{j}({\bf t})=y_{j},\qquad\theta_{j}({\bf t})=\theta_{j}-\sum_{k=1}^{n-1}u^{k}_{j}\,t_{k},\qquad\text{ for }k=2,\ldots,n.$
-From this we have:
-(81) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})Y,$
-where $\alpha_{k}(\mathbf{t})=\exp\left(i(\theta_{k+1}(\mathbf{t})-\theta_{k}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{n+1}(\mathbf{t})=0$ and $Y$ is defined in (33).
-The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
-5.3. The multi-time dynamics on low-dimensional strata.
-When $\psi_{i}=0$ except for $i=j_{0},j_{1},\ldots,j_{s}$ (here $j_{0}=1$), the matrix $g$ can be written as
-(82) $g=\alpha PY.$
-In coordinates $(y,\theta)$ the symplectic form on the correspond stratum of $S(\mathcal{O})$, see remark 3, is
-(83) $\Omega=-\sum_{a=1}^{s}\mathrm{d}y_{j_{a}}\wedge\mathrm{d}\theta_{j_{a}},$
-which implies
-$\{y_{j_{a}},y_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},\theta_{j_{b}}\}=0,\qquad\{\theta_{j_{a}},y_{j_{b}}\}=\delta_{ab}.$
-Hamiltonian vector fields $v_{k}$ restrict to $S(\mathcal{O})$ and have the form
-(84) $v_{k}=\sum_{i=1}^{n}\sum_{a=1}^{s}x_{i}^{k-1}\frac{\partial x_{i}}{\partial y_{j_{a}}}\frac{\partial}{\partial\theta_{j_{a}}},\qquad\text{ for }k=2,\ldots,n.$
-Here we can choose $s$ independent vector fields. In particular, we can choose the first $s$ vectors.
-Here we can express $x_{i}$ in terms of $y_{i}$ according to (36):
-(85) $v_{k}=\sum_{a=2}^{s}u^{k}_{a}\frac{\partial}{\partial\theta_{j_{a}}},$
-where
-(86) $u^{k}_{a}=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\sum_{a=1}^{s}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right)(y_{j_{a}}-c)\right)^{k-1}\left(\frac{n+1-j_{a}}{n}-\delta_{i\geq j_{a}}\right).$
-The solution to the initial value problem $y_{j_{a}}(0)=y_{j_{a}},\,\theta_{j_{a}}(0)=\theta_{j_{a}}$ for these equations is
-(87) $y_{j_{a}}({\bf t})=y_{j_{a}},\qquad\theta_{j_{a}}({\bf t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{k}_{a}\,t_{k},\qquad\text{ for }a=1,\ldots,s.$
-So,
-(88) $x({\bf t})=x,\qquad g({\bf t})=\alpha(\mathbf{t})P\,Y,$
-where $\alpha_{j_{a}}(\mathbf{t})=\exp\left(i(\theta_{j_{a}}(\mathbf{t})-\theta_{j_{a+1}}(\mathbf{t}))\right)$ with $\theta_{1}(\mathbf{t})=\theta_{j_{s+1}}(\mathbf{t})=0$, $Y$ is defined in (48) and $\alpha$ is defined in (50).
-The time dependence of coordinates $q_{j}$ is determined by $g(\mathbf{t})$ since $\{e^{iq_{j}(\mathbf{t})}\}$ are eigenvalues of $g(\mathbf{t})$.
-Example 3.
-Zero-dimensional stratum.
-On this stratum, particles are in equilibrium. Their position are terminated by eigenvalues of (53)
-(89) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{n}=0.$
-Example 4.
-One-dimensional strata.
-There are $(n-1)$ such strata, labeled by $j=2,\ldots,n$, with coordinates $(y_{j},\theta_{j})$ defined in example 2.
-Positions of particles are eigenvalues of $g(t)$ and thus are roots of the characteristic equation
-(90) $\mathrm{det}\left(\lambda-g\right)=\lambda^{n}+(-1)^{j-1}q\beta^{-1}\lambda^{n-j+1}+(-1)^{n-j+1}q\beta\lambda^{j-1}+(-1)^{n}=0,$
-where $\beta=\beta(t)=\alpha_{1}^{-1}=\alpha_{j}=\exp\left(i\theta_{j}(t)\right)=\exp(i\,(\theta_{j}-u\,t))$ and $10<q<$ 1 defined in example 2, and
-(91) $u=\sum_{i=1}^{n}\left(\frac{n-1}{2}c-(i-1)c+\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)(y_{j}-c)\right)\left(\frac{n+1-j}{n}-\delta_{i\geq j}\right)=\\ (n-j+1)(j-1)\left(\frac{c}{2}+\frac{y-c}{n}\right).$
-Appendix A Lemma about determinant of sum of matrix plus rank $1$ matrix
+(103) $u^{F,\{j\}}_{a}(y)=\frac{n+1-j_{a}}{n}\sum_{i=1}^{j_{a}-1}\frac{\partial F}{\partial x_{i}}(x(y))-\sum_{i=j_{a}}^{n}\frac{\partial F}{\partial x_{i}}(x(y)).$
+Here the function $x(y)$ is given by (79).
+As a consequence, the multi-time dynamics in $(y,\theta)$ coordinates is
+(104) $y_{j_{a}}(\mathbf{t})=y_{j_{a}},\ \ \theta_{j_{a}}(\mathbf{t})=\theta_{j_{a}}-\sum_{k=1}^{s}u^{F_{k},\{j\}}_{a}(y)t_{k},$
+where $\mathbf{t}=(t_{1},\dots,t_{s})$. The dynamics is linear in angle variables, as it should be.
+Appendix A Lemma about the determinant of the sum of a matrix plus a rank $1$ matrix
 Lemma 1.
-Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a $1$-dimensional projector. Then
-(92) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{\beta}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z)$
+Let $D$ be a diagonal matrix $D_{ab}=D_{a}\delta_{ab}$ and $P_{ab}=\varphi_{a}\psi_{b}$ be a 11–dimensional projector. Then
+(105) $\det(D-z-P)=\prod_{a}(D_{a}-z)-\sum_{b}\varphi_{b}\psi_{b}\prod_{\begin{subarray}{c}a\neq b\end{subarray}}(D_{a}-z).$
 Proof.
@@
 (107) $=\exp\left(\mathrm{Tr}\ln(D-z)+\operatorname{Tr}\ln\left(1-(D-z)^{-1}P\right)\right)$
-(95) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)$
+(108) $=\prod_{\alpha}(D_{\alpha}-z)\exp\left(\sum_{k\geq 1}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right).$
 Now
 (109) $\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}$ $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{n}}-z)^{-1}P_{a_{1}a_{2}}\cdots P_{a_{k}a_{1}}$
-(97) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}$
+(110) $=\sum_{a_{1},\ldots,a_{k}}(D_{a_{1}}-z)^{-1}\cdots(D_{a_{k}}-z)^{-1}\prod_{i=1}^{k}\varphi_{a_{i}}\psi_{a_{i}}.$
 From here:
-$\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}$
+(111) $\exp\left(\sum_{k=1}^{\infty}\frac{1}{k}\mathrm{Tr}\left((D-z)^{-1}P\right)^{k}\right)=1-\sum_{b=1}^{n}(D_{b}-z)^{-1}\varphi_{b}\psi_{b}.$
 Together with (106) this implies (105). ∎